### v5

# Good involutions of conjugation subquandles

Abstract.
Posed by Taniguchi, the classification of quandles with good involutions is a difficult question with applications to surface-knot theory. We address this question for subquandles of conjugation quandles, including all core quandles. We also study good involutions of faithful racks. In particular, we obtain sharp bounds on the number of good involutions of racks in these families.
As an application of our results, we implement group-theoretic algorithms that compute all good involutions of conjugation quandles and core quandles; we provide data for those up to order 232323. As another application, we construct infinite families of connected, noninvolutory symmetric quandles.
We also classify symmetric and Legendrian racks, quandles, and kei up to order 888 using a computer search. Finally, we exhibit an equivalence of categories between racks and Legendrian racks that induces an equivalence between involutory racks, Legendrian kei, and symmetric kei.
Key words and phrases:
Classification, conjugation, core quandle, enumeration, good involution, group, quandle, rack, symmetric quandle
2020 Mathematics Subject Classification:
Primary 20N02; Secondary 20E34, 20E45, 57K12
1. Introduction
[…]
In this paper, we study the following classification problems posed by Taniguchi [taniguchi].
Problem 1.1.
Find necessary and sufficient conditions for good involutions $\rho$ of quandles to exist.
Problem 1.2.
Given a quandle $R$, classify all symmetric quandles of the form $(R,\rho)$.
[…]
This paper addresses Problems 1.1 and 1.2 for large classes of quandles of broad interest called conjugation quandles and core quandles. These quandles are constructed using the functors $\operatorname{Conj}$ and $\operatorname{Core}$ from the category of groups to the category of quandles. Joyce [joyce] and Matveev [matveev] independently introduced these quandles to generalize Wirtinger presentations of knot groups and develop complete invariants of knots.
[…]
1.1. Overview of results
Our main results are the following. Theorem 1.3 states that good involutions $\rho$ of conjugation subquandles $\operatorname{Conj}X$ of conjugation quandles $\operatorname{Conj}G$ factor through the natural projection $\pi:X\twoheadrightarrow X/\langle X\rangle$ and a unique function $\zeta^{*}:X/\langle X\rangle\to Z(\langle X\rangle)$, where $X/\langle X\rangle$ denotes the set of orbits of $X$ under the action of the subgroup $\langle X\rangle\leq G$ by conjugation. Moreover, $\zeta^{*}$ completely determines $\rho$. Theorem 1.4 gives an analogous statement for core quandles $\operatorname{Core}G$.
Theorem 1.3.
Let $G$ be a group, and let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G=(G,s)$. Then for all functions $\rho:X\to X$, $\rho$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a function $\zeta:X\to Z(\langle X\rangle)$ such that
$\rho(x)=(\zeta(x)x)^{-1},\qquad\zeta s_{x}=\zeta=\zeta\rho$
for all $x\in X$. (See Theorem 6.7; cf. Remark 6.9 and Corollary 6.11.)
Theorem 1.4.
Let $G$ be a group such that the core quandle $\operatorname{Core}G=(G,s)$ is nontrivial, and let $T$ be the subgroup
$T:=\{z\in Z(G)\mid z^{2}=1\}$
of the center $Z(G)$. Then for all functions $\rho:G\to G$, $\rho$ is a good involution of $\operatorname{Core}G$ if and only if there exists a function $\alpha:G\to T$ such that
$\rho(g)=\alpha(g)g,\qquad\alpha s_{g}=\alpha=\alpha\rho$
for all $g\in G$. (See Theorem 7.11; cf. Remark 7.13 and Corollaries 7.16 and 7.17.)
For example, it is well-known that if $\operatorname{Conj}X$ is a subquandle of $\operatorname{Conj}G$ and $X$ is closed under inverses, then the inversion map $x\mapsto x^{-1}$ is a good involution of $\operatorname{Conj}X$. This good involution corresponds to letting $\zeta:X\to Z(\langle X\rangle)$ be the constant function $x\mapsto 1$.
As applications of Theorem 1.3 and 1.4, we describe procedures (Algorithms 8.1 and 8.5) for constructing all good involutions of conjugation subquandles and core quandles, and we provide data for groups up to order 232323 in a GitHub repository [code]. This addresses Problem 1.2.
1.1.1. Computer search
Table 1. Enumeration of isomorphism classes of various types of symmetric racks up to order 888.
Order 0 1 2 3 4 5 6 7 8
Symmetric racks 1 1 4 9 42 154 1064 6678 73780
Symmetric quandles 1 1 2 5 13 44 187 937 6459
Symmetric kei 1 1 2 5 13 42 180 906 6317
Through an exhaustive search in the computer algebra system GAP [GAP4], we tabulated all isomorphism classes of symmetric racks, quandles, and kei up to order 888. This answers Problem 1.2 for all racks up to order 888. Our classification used Vojtěchovský and Yang’s [library] library of racks in a way similar to the algorithms described in [taGL]*App. A.
We provide our code and data in a GitHub repository [code]. Table 1 enumerates our data.
1.1.2. CNS-quandles
Connected, noninvolutory symmetric quandles, which we call CNS-quandles, are particularly helpful for studying 333-manifolds [nosaka] and nonorientable surface-knots [dihedral]. However, CNS-quandles are difficult to construct. For example, generalized Alexander quandles have good involutions if and only if they are kei [taniguchi]*Thm. 1.3, and our computer search showed the following.
Proposition 1.5.
Up to isomorphism, the only CNS-quandle of order 888 or lower is the symmetric Galkin quandle of order 666 constructed by Clark et al. [galkin-symm]*Prop. 5.13.
Thus, the following problems arise; see [galkin-symm]*Sec. 5.
Problem 1.6.
Construct infinite families of CNS-quandles.
Problem 1.7.
For which integers $k\in\mathbb{Z}^{+}$ do there exist CNS-quandles of order $k$?
[…]
Proposition 1.8.
For all integers $m\geq 3$ and powers $q$ of odd primes, and for all integers $n\geq 5$ and $k$ such that $3\leq k\leq 2\lfloor n/2\rfloor$, there exist CNS-quandles of orders
$\frac{(q^{m}-1)(q^{m-1}-1)}{q-1},\qquad\frac{n!}{k(n-k)!}.$
(See Examples 9.3 and 9.4 for details.)
Remark 1.9.
If $q=3=k$, then for all integers $m\geq 3$ and $n\geq 5$ such that $n\not\equiv 0,1,2$ mod 999, the CNS-quandles in Proposition 1.8 do not share orders with the CNS-quandles constructed by Clark et al. Most of them also do not share orders with the CNS-quandles constructed by Carter et al.
1.1.3. Connections with Legendrian racks
[…]
Generalizing prior work of Kulkarni and Prathamesh [original] in 2017, Ceniceros et al. [ceniceros] introduced Legendrian racks in 2021. In 2023, Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced generalizations of Legendrian racks called GL-racks. In 2025, the author [taGL] studied algebraic aspects of GL-racks and classified them up to order 888, and Cheng and He [he] showed that GL-racks distinguish Legendrian knots up to the absolute values of their classical invariants.
In this paper, we show that the classification problems for symmetric kei, Legendrian kei, and involutory racks are equivalent. We also obtain similar results for other types of GL-racks.
Theorem 1.10.
The categories of symmetric kei, Legendrian kei, and involutory racks are equivalent. In fact, good involutions and Legendrian structures coincide for kei. (See Theorem 10.1.)
In particular, the bottom row of Table 1 enumerates isomorphism classes of Legendrian kei and those of involutory racks; cf. Table 5 in Section 10.
1.2. Structure of the paper
In Section 2, we provide definitions and examples of racks, quandles, and kei. In Section 3, we do the same for rack antiautomorphisms and good involutions.
In Section 4, we describe good involutions in terms of rack (anti)automorphisms. In particular, we deduce that only self-dual racks have good involutions. This addresses Problem 1.1.
In Section 5, we discuss Problems 1.1 and 1.2 for faithful racks. A class of examples shows that infinitely many quandles have involutory antiautomorphisms but no good involutions.
In Section 6, we prove Theorem 1.3 and discuss some special cases along the way. This addresses Problems 1.1 and 1.2 for subquandles of conjugation quandles.
In Section 7, we prove Theorem 1.4 by reducing to a special case of Theorem 1.3. This addresses Problems 1.1 and 1.2 for core quandles.
In Section 8, we describe procedures for computing all good involutions of conjugation subquandles and core quandles. We provide data for groups up to order 232323 [code]. This addresses Problem 1.2.
In Section 9, we compute all good involutions of several classes of connected quandles. Along the way, we obtain Proposition 1.8. This addresses Problems 1.6 and 1.7.
In Section 10, we construct an equivalence of categories between racks and Legendrian racks. We discuss various restrictions of this equivalence; in particular, we obtain Theorem 1.10.
In Section 11, we pose several questions for further research.
Notation.
We use the following notation throughout this paper. Denote the composition of functions $\varphi:X\to Y$ and $\psi:Y\to Z$ by $\psi\varphi$. Given a set $X$, let $\operatorname{id}_{X}$ denote the identity map of $X$, and let $S_{X}$ denote the symmetric group of $X$. Given an integer $n\in\mathbb{Z}^{+}$, let $S_{n}$ and $A_{n}$ be the symmetric and alternating groups on $n$ letters, respectively.
Acknowledgments
I thank Samantha Pezzimenti, Jose Ceniceros, and Peyton Wood for respectively introducing me to knot theory, quandles, and Legendrian racks. I also thank Sam Raskin for advising me during the writing of [taGL], the results of which helped inspire Theorem 1.10.
2. Preliminaries: Racks
2.1. Racks and quandles
Definition 2.1.
[…]
$s_{x}s_{y}=s_{s_{x}(y)}s_{x}$
[…]
Remark 2.2.
[…]
$s_{y}(x)=x\triangleright y.$
Remark 2.3.
Although the empty set can be viewed as a rack, we will only discuss nonempty racks in this paper.
Definition 2.4.
Let $R:=(X,s)$ be a rack.
  • We say that $R$ is a quandle if $s_{x}(x)=x$ for all $x\in X$.
  • If $R$ is a quandle and $Y\subseteq X$ is a subset such that $s_{y}(Y)=Y$ for all $y\in Y$, then we say that the quandle $(Y,s|_{Y})$ is a subquandle of $R$.
  • We say that $R$ is involutory if $s_{x}^{2}=\operatorname{id}_{X}$ for all $x\in X$.
  • A kei is an involutory quandle.
  • We say that $R$ is faithful if $s$ is injective.
Example 2.5 ([fenn]).
Let $X$ be a set, and fix $\sigma\in S_{X}$. Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}=\sigma$ for all $x\in X$. Then $R:=(X,s)$ is a rack called a permutation rack, cyclic rack, or constant action rack.
Note that $R$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$, in which case we call $R$ a trivial quandle. Moreover, $R$ is involutory if and only if $\sigma$ is an involution, and $R$ is faithful if and only if $|X|\leq 1$.
[…]
Let $G$ be a group, and define $s:G\rightarrow S_{G}$ by sending each element $g\in G$ to the conjugation map defined by
$s_{g}(h):=ghg^{-1}$
for all $h\in G$. Then $\operatorname{Conj}G:=(G,s)$ is a quandle called a conjugation quandle or conjugacy quandle. Note that $s_{g}^{-1}=s_{g^{-1}}$ for all $g\in G$.
[…]
Let $G$ be a group, and define $s:G\to S_{G}$ by
$s_{g}(h):=gh^{-1}g$
for all $g,h\in G$. Then $\operatorname{Core}G:=(G,s)$ is a kei called the core quandle of $G$.
Note that $\operatorname{Core}G$ is faithful if and only if the center $Z(G)$ is 222-torsionless [core2]*Lem. 4.1. Moreover, $\operatorname{Core}G$ is a permutation rack if and only if $G$ is an elementary abelian 222-group, in which case $\operatorname{Core}G$ is a trivial quandle.
Furthermore, if $G$ is an additive abelian group, then the core quandle $T(G):=\operatorname{Core}(G)$ is called a Takasaki kei [quandlebook]*Ex. 54. Note that $T(G)$ is faithful if and only if $G$ is 222-torsionless.
In particular, if $G$ is cyclic of order $n$, then we call $R_{n}:=T(\mathbb{Z}/n\mathbb{Z})$ a dihedral quandle. By the above discussion, $R_{n}$ is faithful if and only if $n$ is odd.
Example 2.8 ([taniguchi], cf. [alex]).
[…]
$s_{g}(h):=\varphi(hg^{-1})g$
for all $g,h\in G$. A routine check shows that $\operatorname{GAlex}(G,\varphi)$ is faithful if and only if the identity element of $G$ is the only fixed point of $\varphi$.
In particular, let $G$ be an additive abelian group, and let $\varphi$ be the inversion map $g\mapsto-g$. Then $\operatorname{GAlex}(G,\varphi)$ equals the Takasaki kei $T(G)$.
Example 2.9 ([tetrahedral]*Ex. 2.8).
Define the set $X:=\{1,2,3,4\}$, and define $s:X\to S_{4}$ by
$s_{1}:=(234),\quad s_{2}:=(143),\quad s_{3}:=(124),\quad s_{4}:=(132)$
in cycle notation. Then $(X,s)$ is a faithful, noninvolutory quandle called the tetrahedral quandle or regular tetrahedron quandle.
2.1.1. Conjugation subquandles
We discuss subquandles of conjugation quandles. In the following, let $G$ be a group, and let $s:G\to S_{G}$ denote the rack structure of $\operatorname{Conj}G$.
Lemma 2.10 ([conjugation]*Thm. 3.1).
For all subsets $X\subseteq G$, the pair $\operatorname{Conj}X:=(X,s|_{X})$ is a subquandle of $\operatorname{Conj}G$ if and only if $X$ is closed under conjugation by elements of the subgroup $\langle X\rangle$ of $G$ generated by $X$.
Observation 2.11.
Let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$. Then $\operatorname{Conj}X$ is a kei if and only if $x^{2}\in Z(\langle X\rangle)$ for all $x\in X$, and $\operatorname{Conj}X$ is faithful if and only if the action of $\langle X\rangle$ on $X$ by conjugation is faithful. In particular, $\operatorname{Conj}G$ is faithful if and only if $G$ is centerless.
Example 2.12.
The conjugation quandle of the quaternion group $Q_{8}$ is a nonfaithful kei.
Of particular interest in this paper are subquandles $\operatorname{Conj}X$ of conjugation quandles $\operatorname{Conj}G$ whose underlying sets $X$ are symmetric subsets of $G$.
Definition 2.13.
Let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$. We say that $\operatorname{Conj}X$ is inversion-closed [^1] We chose this name over the name “symmetric subquandle” to avoid confusion with Definition 3.8. if $x^{-1}\in X$ for all $x\in X$.
Remark 2.14.
By Lemma 2.10, a subquandle $\operatorname{Conj}X$ of $\operatorname{Conj}G$ is inversion-closed if and only if every element $x\in X$ is conjugate to its inverse $x^{-1}$ in the subgroup $\langle X\rangle\leq G$.
Example 2.15.
Let $n\in\mathbb{Z}^{+}$ be a positive integer, and let $X$ be any conjugacy class in the symmetric group $S_{n}$. Then $\operatorname{Conj}X$ is an inversion-closed subquandle of $\operatorname{Conj}S_{n}$.
Example 2.16.
Let $X$ be the split conjugacy class
$X:=\{(234),(143),(124),(132)\}$
in the alternating group $A_{4}$. Then $\langle X\rangle=A_{4}$ is centerless, so $\operatorname{Conj}X$ is a faithful, noninvolutory subquandle of $\operatorname{Conj}A_{4}$. Moreover, $\operatorname{Conj}X$ is not inversion-closed.
2.1.2. Rack homomorphisms
[…]
Definition 2.17.
[…]
$\varphi s_{x}=t_{\varphi(x)}\varphi$
for all $x\in X$. A rack isomorphism is a bijective rack homomorphism, and a rack automorphism of is an isomorphism from a rack to itself.
[…]
It is straightforward to check that the assignments $G\mapsto\operatorname{Conj}G$ and $G\mapsto\operatorname{Core}G$ define faithful functors from the category of groups to the category of quandles.
Example 2.19.
The tetrahedral quandle from Example 2.9 is isomorphic to the quandle in Example 2.16.
Definition 2.20.
We denote the automorphism group of a rack $R=(X,s)$ by $\operatorname{Aut}R$. The inner automorphism group [^2] Some authors call $\operatorname{Inn}R$ the right multiplication group $\operatorname{RMlt}R$ or operator group $\operatorname{Op}R$ of $R$. $\operatorname{Inn}R$ of $R$ is the subgroup of $S_{X}$ generated by $S(X)$:
$\operatorname{Inn}R:=\langle s_{x}\mid x\in X\rangle.$
Observation 2.21.
For all racks $R$, the group $\operatorname{Inn}R$ is a normal subgroup of $\operatorname{Aut}R$.
Example 2.22 ([fenn]*Sec. 1.3).
Let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle$ be the subgroup of $G$ generated by $X$. Then $\operatorname{Inn}(\operatorname{Conj}X)$ equals the inner automorphism group $\operatorname{Inn}H$ of the group $H$. In particular, the assignment $s_{x}\mapsto xZ(H)$ is a group isomorphism from $\operatorname{Inn}(\operatorname{Conj}X)$ to $H/Z(H)$.
Example 2.23.
The inner automorphism group of the tetrahedral quandle from Example 2.9 is the alternating group $A_{4}$.
Definition 2.24.
We say that a rack $R$ is connected if the action of $\operatorname{Inn}R$ on $R$ is transitive.
Lemma 2.25 ([conjugation]*Thm. 3.2).
If $\operatorname{Conj}X$ is a subquandle of $\operatorname{Conj}G$, then $\operatorname{Conj}X$ is connected if and only if $X$ is a conjugacy class in the subgroup $\langle X\rangle\leq G$.
Lemma 2.26 ([core2]*Cor. 3.7).
Let $G$ be a group. Then the core quandle $\operatorname{Core}G$ is connected if and only if $G$ is generated by the squares of its elements.
3. Preliminaries: Good involutions
3.1. Rack antiautomorphisms
A key ingredient in our proofs is rack antiautomorphisms, which Horvat [antiaut] introduced in 2017 to study homeomorphisms of pairs $(S^{3},K)$ for all knots $K\subset S^{3}$.
Definition 3.1.
Given a rack $R=(X,s)$, define $s^{-1}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$. Then the rack $R^{\mathrm{op}}:=(X,s^{-1})$ is a rack called the dual rack of $R$.
Definition 3.2.
Given a rack $R$, an antiautomorphism of $R$ is a rack isomorphism $\rho:R\xrightarrow{\sim}R^{\mathrm{op}}$. Let $\operatorname{Anti}R$ be the set of antiautomorphisms of $R$. We say that $R$ is self-dual if $\operatorname{Anti}R$ is nonempty.
Concretely, $\operatorname{Anti}R$ is the set of permutations $\rho\in S_{X}$ that satisfy
$\rho s_{x}=s^{-1}_{\rho(x)}\rho$
for all $x\in X$.
Observation 3.3 ([antiaut]*Rem. 2.6).
The composition of two rack antiautomorphisms is a rack automorphism, while the composition of an antiautomorphism and an automorphism (in either order) is an antiautomorphism.
Example 3.4.
Let $R=(X,s)$ be a rack. Then the following are equivalent:
  • $R$ is involutory.
  • $R$ is equal (not only isomorphic) to the dual rack $R^{\mathrm{op}}$.
  • $\operatorname{id}_{X}\in\operatorname{Anti}R$.
Under the above conditions, $R$ is self-dual, and $\operatorname{Anti}R=\operatorname{Aut}R$.
Example 3.5.
Let $G$ be a group, let $\iota\in S_{G}$ be the inversion map $g\mapsto g^{-1}$, and let $\operatorname{Conj}X$ be an inversion-closed subquandle of $\operatorname{Conj}G$. Then the restriction $\iota|_{X}$ is an involution of $X$. Moreover, $\iota|_{X}$ is an antiautomorphism of $\operatorname{Conj}X$ because
$\iota s_{y}(x)=(yxy^{-1})^{-1}=yx^{-1}y^{-1}=\iota(y)^{-1}\iota(x)\iota(y)=s^{-1}_{\iota(y)}\iota(x)$
for all $x,y\in X$. In particular, $\operatorname{Conj}X$ is self-dual. Note that $\iota\in\operatorname{Aut}(\operatorname{Conj}X)$ if and only if $\operatorname{Conj}X$ is a kei.
Example 3.6 ([tetrahedral]*Ex. 2.16).
The tetrahedral quandle $Q$ from Example 2.9 is self-dual; indeed, $\operatorname{Anti}Q$ contains every transposition in $S_{4}$.
Example 3.7.
Not all quandles are self-dual. For example, let $n\in\mathbb{Z}^{+}$ and $a\in\mathbb{Z}$ be relatively prime integers, define the quotient ring $X:=(\mathbb{Z}/n\mathbb{Z})[t^{\pm 1}]/(t-a)$, and define $\varphi:X\to X$ by $x\mapsto(1-t)x$. Then $\operatorname{GAlex}(X,\varphi)$ is called a linear Alexander quandle. A result of Nelson [alex]*Cor. 2.8 states that $\operatorname{GAlex}(X,\varphi)$ is self-dual if and only if $a$ is a square modulo $n/{\gcd(n,1-a)}$.
3.2. Good involutions
This paper concerns the category of symmetric racks, which Kamada [symm-quandles] introduced in 2007.
Definition 3.8.
Let $R=(X,s)$ be a rack. A good involution of $R$ is an involution $\rho\in S_{X}$ that satisfies the equalities
$\rho s_{x}=s_{x}\rho,\qquad s_{\rho(x)}=s^{-1}_{x}$
for all $x\in X$. We denote the set of good involutions of $R$ by $\operatorname{Good}R$.
A symmetric rack (resp. symmetric quandle) is a pair $(R,\rho)$ where $R$ is a rack (resp. quandle) and $\rho\in\operatorname{Good}R$.
Observation 3.9.
The set of good involutions is a rack invariant. Explicitly, if $\varphi:R_{1}\xrightarrow{\sim}R_{2}$ is a rack isomorphism, then the assignment $\rho\mapsto\varphi\rho\varphi^{-1}$ is a bijection from $\operatorname{Good}R_{1}$ to $\operatorname{Good}R_{2}$.
Symmetric racks form a category with the following morphisms; symmetric quandles form a full subcategory of this category.
Definition 3.10.
A symmetric rack homomorphism between two symmetric racks $(R_{1},\rho_{1})$ and $(R_{2},\rho_{2})$ is a rack homomorphism $\varphi$ from $R_{1}$ to $R_{2}$ that intertwines the good involutions:
$\varphi\rho_{1}=\rho_{2}\varphi.$
A symmetric rack isomorphism is a bijective symmetric rack homomorphism.
Example 3.11 ([symm-quandles-2]*Prop. 3.1).
Given a quandle $R=(X,s)$, let $I\subset S_{X}$ the set of all involutions of $X$. Then $\operatorname{Good}R=I$ if and only if $R$ is a trivial quandle.
Example 3.12 ([symm-quandles]*Ex. 2.4).
Let $G$ be a group. Then the inversion map $\iota\in S_{G}$ defined by $g\mapsto g^{-1}$ is a good involution of $\operatorname{Conj}G$.
Example 3.13.
Example 3.12 immediately implies its own generalization. Let $G$ be a group, let $\iota\in S_{G}$ be the inversion map $g\mapsto g^{-1}$, and let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$. Then the restriction $\iota|_{X}$ is a good involution of $\operatorname{Conj}X$ if and only if $\operatorname{Conj}X$ is inversion-closed; cf. Section 6.
Example 3.14 ([karmakar]*Ex. 2.3).
Let $G$ be a group, and let $z\in Z(G)$ be a central element of order 111 or 222. Then multiplication by $z$ is a good involution of $\operatorname{Core}G$.
Example 3.15.
Let $R=\operatorname{GAlex}(G,\varphi)$ be a generalized Alexander quandle. A theorem of Taniguchi [taniguchi]*Thm. 1.3 states that $\operatorname{Good}R$ is nonempty if and only if $R$ is a kei.
3.2.1. Involutory symmetric racks
We recall a well-known result of Kamada and Oshiro [symm-quandles-2].
Proposition 3.16 ([symm-quandles-2]*Prop. 3.4).
Let $R=(X,s)$ be a rack. Then the following statements are equivalent:
  • $R$ is involutory.
  • $\operatorname{id}_{X}\in\operatorname{Good}R$.
  • $\operatorname{Good}R$ contains an element of $\operatorname{Aut}R$.
  • $\operatorname{Good}R$ is a nonempty subset of $\operatorname{Aut}R$.
Remark 3.17.
Although Kamada and Oshiro only stated their result for quandles, their proof holds more generally for racks.
In fact, we have the following.
Proposition 3.18.
If $R$ is involutory, then $\operatorname{Good}R$ is precisely the set $I$ of involutions $\rho:X\to X$ contained in the centralizer $C_{\operatorname{Aut}R}(\operatorname{Inn}R)$.
Proof.
The containment $\operatorname{Good}R\subseteq I$ follows from Proposition 3.16. Conversely, if $\rho\in I$, then
$s_{\rho(x)}=\rho s_{\rho(x)}\rho=\rho^{2}s_{x}=s_{x}=s_{x}^{-1},$
so $\rho$ is a good involution of $R$. ∎
Corollary 3.19.
Suppose that $R$ is involutory. If $\operatorname{Aut}R$ is a simple group, then $\operatorname{Good}R$ is either $\{\operatorname{id}_{X}\}$ or the set of all involutory automorphisms of $R$.
Proof.
Since $\operatorname{Inn}R$ is a normal subgroup of $\operatorname{Aut}R$, so is the centralizer $C_{\operatorname{Aut}R}(\operatorname{Inn}R)$. Hence, the claim follows from Proposition 3.18. ∎
4. General results for racks
Let $R=(X,s)$ be a rack. In this section, we characterize good involutions $\rho\in\operatorname{Good}R$ as involutory central antiautomorphisms of $R$ (Proposition 4.2). As a result, self-duality is a necessary condition for $R$ to have good involutions (Corollary 4.4), and every good involution factors through some automorphism $\varphi$ and a given antiautomorphism $\iota$ (Corollary 4.7).
4.1. Good involutions are antiautomorphisms
We record a “two-implies-three” rule for good involutions. This will facilitate the proof of Theorem 1.3 later in the paper.
Lemma 4.1.
For all permutations $\rho\in S_{X}$, any pair of the following three conditions implies the other:
[…]
    For all $x\in X$, $s_{\rho(x)}^{-1}=s_{x}$.
[…]
    $\rho$ lies in the centralizer $C_{S_{X}}(\operatorname{Inn}R)$.
[…]
    $\rho$ is an antiautomorphism of $R$.
Proof.
Straightforward. ∎
Proposition 4.2.
For all involutions $\rho\in S_{X}$, the following are equivalent:
  • $\rho$ is a good involution of $R$.
  • $\rho$ satisfies at least two out of the three conditions in Lemma 4.1.
  • $\rho$ satisfies all three of the conditions in Lemma 4.1.
Proof.
Combined with the assumption that $\rho$ is an involution, conditions 1 and 2 form the definition of a good involution. Hence, the claim follows immediately from Lemma 4.1. ∎
Remark 4.3.
Together, Example 3.4 and Proposition 4.2 recover Proposition 3.16.
The following addresses Problem 1.2.
Corollary 4.4.
If $R$ is not self-dual, then $R$ has no good involutions.
Proof.
By definition, $R$ is self-dual if and only if $\operatorname{Anti}R$ is nonempty. By Proposition 4.2, $\operatorname{Good}R$ is a subset of $\operatorname{Anti}R$, so the claim follows. ∎
Remark 4.5.
The inverse of Corollary 4.4 is not true in general, even if $\operatorname{Anti}R$ contains involutions. For example, all transpositions in the symmetric group $S_{4}$ are antiautomorphisms of the tetrahedral quandle $Q$ from Example 2.9. Nevertheless, Table 1 implies that $Q$ has no good involutions; cf. Example 6.6. In fact, there are infinitely many examples of this nature; see Proposition 5.10.
4.2. Factoring good involutions
In light of Corollary 4.4, assume for the rest of this section that $R$ is self-dual, and fix an antiautomorphism $\iota\in\operatorname{Anti}R$.
Proposition 4.6.
The assignment $\varphi\mapsto\iota\varphi$ is a bijection from $\operatorname{Aut}R$ to $\operatorname{Anti}R$.
Proof.
Since $R\cong R^{\mathrm{op}}$, this follows from Observation 3.3 and the definition of $\operatorname{Anti}R$. ∎
The following corollaries follow directly from Propositions 4.2 and 4.6. In particular, Corollary 4.7 states that knowing the elements of $\operatorname{Aut}R$ is sufficient for computing $\operatorname{Good}R$.
Corollary 4.7.
For all involutions $\rho\in S_{X}$, $\rho$ is a good involution of $R$ if and only if there exists an automorphism $\varphi\in\operatorname{Aut}R$ such that $\rho=\iota\varphi$ and at least one of the conditions 1, 2 hold.
Corollary 4.8.
The number of good involutions of $R$ is at most the order of $\operatorname{Aut}R$:
$|{\operatorname{Good}R}|\leq|{\operatorname{Anti}R}|=|{\operatorname{Aut}R}|.$
Equality holds if and only if $\operatorname{Anti}R$ only contains involutions in the centralizer $C_{S_{X}}(\operatorname{Inn}R)$.
Remark 4.9.
The bound in Corollary 4.8 is sharp in the sense that there exist quandles $R$ of infinitely many orders $n$ such that ${\operatorname{Good}R}={\operatorname{Anti}R}$. To see this, let $n\geq 2$ be an even integer, and define a rack structure $s$ on the set $X=\{1,2,\dots,n\}$ in cycle notation by
$s_{1},s_{2}:=\operatorname{id}_{X};\qquad s_{k},s_{k+1}:=(k-2\;\;k-1)$
for all odd $3\leq k<n$. Then $R:=(X,s)$ is a kei, so Example 3.4 implies that $\operatorname{Anti}R$ is the elementary abelian 222-group
$\operatorname{Anti}R=\operatorname{Aut}R=\langle(12),(34),\dots,(n-1\;\;n)\rangle\cong(\mathbb{Z}/2\mathbb{Z})^{n/2}.$
In particular, $R$ satisfies the last condition in Corollary 4.8, so ${\operatorname{Good}R}={\operatorname{Anti}R}$.
5. Faithful symmetric racks
In this section, we study Problems 1.1 and 1.2 for faithful racks.
5.1. General results
Proposition 5.1.
If $R$ is faithful, then $R$ has at most one good involution.
Proof.
Let $\rho,\rho^{\prime}\in\operatorname{Good}R$. For all $x\in X$,
$s_{\rho(x)}=s^{-1}_{x}=s_{\rho^{\prime}(x)},$
so $\rho(x)=\rho^{\prime}(x)$ because $R$ is faithful. Therefore, $\rho=\rho^{\prime}$. ∎
Remark 5.2.
Example 2.9 and Proposition 5.10 give examples of faithful, self-dual quandles with no good involutions.
On the other hand, Clark et al. [galkin-symm]*Prop. 5.13 provided infinitely many faithful, self-dual quandles that do have a (necessarily unique) good involution. The next two results do the same.
Corollary 5.3.
Let $G$ be a group, let $\operatorname{Conj}X$ be an inversion-closed subquandle of $\operatorname{Conj}G$, and let $\langle X\rangle$ be the subgroup of $G$ generated by $X$. If $\langle X\rangle$ is centerless, then $\operatorname{Good}(\operatorname{Conj}X)=\{\iota|_{X}\}$.
Proof.
By assumption, the action of $\langle X\rangle$ on $X$ by conjugation is faithful, so Observation 2.11 states that $\operatorname{Conj}X$ is faithful. Hence, the claim follows from Example 3.13 and Proposition 5.1. ∎
Remark 5.4.
An inverse to Corollary 5.3 holds; see Corollary 6.5.
Corollary 5.5.
If $R$ is faithful and involutory, then $\operatorname{Good}R=\{\operatorname{id}_{X}\}$.
Proof.
This follows from Propositions 3.16 and 5.1. ∎
Example 5.6.
Let $G$ be a group, and let $\varphi\in\operatorname{Aut}G$ be an automorphism of $G$. By a theorem of Taniguchi [taniguchi]*Thm. 1.3, the generalized Alexander quandle $\operatorname{GAlex}(G,\varphi)$ has a good involution if and only if it is a kei. In this case, if the only fixed point of $\varphi$ is the identity element $1\in G$, then Example 2.8 and Corollary 5.5 imply that $\operatorname{Good}(\operatorname{GAlex}(G,\varphi))=\{\operatorname{id}_{G}\}$.
In particular, if $G$ is abelian and 222-torsionless, then the only good involution of the Takasaki kei $T(G)$ is the identity map $\operatorname{id}_{G}$; cf. Corollary 7.17. This generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles $R_{n}$ of odd order.
5.2. Symmetric Galkin quandles
[…]
In particular, Clark et al. [galkin-symm]*Prop. 5.13 constructed good involutions of certain Galkin quandles. We show that no other good involutions of Galkin quandles exist. Aside from addressing Problems 1.1 and 1.2, this is also motivated by finding infinitely many quandles that have involutory antiautomorphisms and no good involutions; cf. Remark 4.5.
Throughout this section, let $A$ be an additive abelian group.
Definition 5.7 ([galkin-1]).
Fix an element $c\in A$. Define functions ${\mu:\mathbb{Z}/3\mathbb{Z}\to\mathbb{Z}}$ and $\tau:\mathbb{Z}/3\mathbb{Z}\to A$ by
$\mu(n):=\begin{cases}2&\text{if }n=0,\\ -1&\text{if }n=1,2,\end{cases}\qquad\tau(n):=\begin{cases}0&\text{if }n=0,1,\\ c&\text{if }n=2.\end{cases}$
Define a rack structure $s$ on the direct product $\mathbb{Z}/3\mathbb{Z}\times A$ by
$s_{(y,b)}(x,a):=(2y-x,-a+\mu(x-y)b+\tau(x-y)).$
Then the pair $\mathcal{G}(A,c):=(\mathbb{Z}/3\mathbb{Z}\times A,s)$ is a quandle called a Galkin quandle.
Remark 5.8.
A similar definition of Galkin quandles also exists [galkin-symm]*Def. 3.4; its equivalence to the above definition was proven in [galkin-symm]*Lem. 3.6.
Remark 5.9 ([galkin-symm]*Lem. 5.2, Prop. 5.11, Cor. 5.12).
a
  • All Galkin quandles are faithful.
  • For all Galkin quandles $\mathcal{G}(A,c)$, the involution $(x,a)\mapsto(-x,a)$ is an antiautomorphism of $\mathcal{G}(A,c)$. In particular, all Galkin quandles are self-dual.
  • $\mathcal{G}(A,c)$ is involutory if and only if $c=0$.
5.2.1. Results
[…]
$\rho_{c}(x,a):=(x,a+c)$
for all $(x,a)\in\mathbb{Z}/3\mathbb{Z}\times A$. Then $\rho_{c}$ is a good involution of $\mathcal{G}(A,c)$.
Proposition 5.10.
If $\mathcal{G}(A,c)$ is a Galkin quandle, then
$\operatorname{Good}(\mathcal{G}(A,c))=\begin{cases}\{\rho_{c}\}&\text{if }2c=0,\\ \emptyset&\text{otherwise.}\end{cases}$
Proof.
If $2c=0$, then $\rho_{c}\in\operatorname{Good}(\mathcal{G}(A,c))$ by the construction of Clark et al. Conversely, for all good involutions $\rho=(\rho_{1},\rho_{2})$ of $\mathcal{G}(A,c)$ and for all $(x,a),(y,b)\in\mathbb{Z}/3\mathbb{Z}\times A$, the expression
$s_{\rho(y,b)}(x,a)=(2\rho_{1}(y)-x,-a+\mu(x-\rho_{1}(y))\rho_{2}(b)+\tau(x-\rho_{1}(y))$
equals the expression
$s_{(y,b)}^{-1}(x,a)=(2y-x,-a+\mu(y-x)b+\tau(y-x)).$
In particular, $\rho_{1}$ must be the identity map $\operatorname{id}_{\mathbb{Z}/3\mathbb{Z}}$. Letting $z:=x-y$, we obtain
$\mu(z)\rho_{2}(b)+\tau(z)=\mu(-z)b+\tau(-z).$
Taking $z$ to be 111 and 222 respectively yields
$-\rho_{2}(b)=-b+c,\qquad-\rho_{2}(b)+c=-b,$
from which we deduce that $2c=0$ and $\rho_{2}(b)=b+c$ for all $b\in A$. In particular, $\rho=\rho_{c}$. ∎
Corollary 5.11.
If $A$ is 222-torsionless, then for all $c\in A$, there are no good involutions of $\mathcal{G}(A,c)$.
6. Symmetric conjugation quandles
In this section, we prove Theorem 1.3 (Theorem 6.7) and two corollaries. We also consider some special cases along the way (Proposition 6.4, Corollary 6.5). Henceforth, let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle$ be the subgroup of $G$ generated by $X$.
6.1. Preliminary results
To prove Theorem 1.3, we use the following lemmas.
Lemma 6.1.
For all involutions $\rho\in S_{X}$, $\rho$ is a good involution of $\operatorname{Conj}X$ if and only if $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$ and $\rho(x)x\in Z(H)$ for all $x\in X$.
Proof.
$\implies$” If $\rho\in\operatorname{Good}(\operatorname{Conj}X)$, then $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$ by Proposition 4.2. We have to show that $\rho(x)xy=y\rho(x)x$ for all $x,y\in X$. Since $\rho$ is a good involution,
$xyx^{-1}=s_{x}(y)=s^{-1}_{\rho(x)}(y)=\rho(x)^{-1}y\rho(x),$
and the desired equality follows.
[…]
$\rho(x)^{-1}Z(H)=xZ(H)$
holds in the quotient group $H/Z(H)$. It follows from Example 2.22 that
$s_{\rho(x)^{-1}}=s_{x}.$
By the definition of $\operatorname{Conj}X$, the left-hand side equals $s_{\rho(x)}^{-1}$. This completes the proof. ∎
Observation 6.2.
An antiautomorphism $\rho$ of $\operatorname{Conj}X$ is a precisely a permutation of $X$ satisfying
$\rho(yxy^{-1})=\rho(y)^{-1}\rho(x)\rho(y)$
for all $x,y\in X$.
Lemma 6.3.
Let $\zeta:X\to Z(H)$ be a function, and suppose that the function $\rho:X\to H$ defined by
$\rho(x):=\ (\zeta(x)x)^{-1}$
is a permutation of $X$. Then $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$ if and only if $\zeta(yxy^{-1})=\zeta(x)$ for all $x,y\in X$.
Proof.
$\implies$” Suppose that $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$. Since $\zeta(X)\subseteq Z(H)$, Observation 6.2 yields
$yxy^{-1}\zeta(yxy^{-1})$ $=\rho(yxy^{-1})^{-1}$
$=\rho(y)^{-1}\rho(x)^{-1}\rho(y)$
$=yxy^{-1}\zeta(y)\zeta(x)\zeta(y)^{-1}$
$=yxy^{-1}\zeta(x)$
for all $x,y\in X$. Hence, $\zeta(yxy^{-1})=\zeta(x)$.
[…]
6.2. Intermediate results
The following results address Problems 1.1 and 1.2.
Proposition 6.4.
If there exists an element $x\in X$ such that $yx\notin Z(H)$ for all $y\in X$, then $\operatorname{Conj}X$ has no good involutions.
Proof.
For all good involutions $\rho$ of $\operatorname{Conj}X$, Lemma 6.1 states that $\rho(x)x\in Z(H)$ for all $x\in X$, so the claim follows. ∎
Corollary 6.5.
Let $\iota:G\to G$ be the inversion map $g\mapsto g^{-1}$. If $H$ is centerless, then
$\operatorname{Good}(\operatorname{Conj}X)=\begin{cases}\{\iota|_{X}\}&\text{if }\iota(X)=X,\\ \emptyset&\text{otherwise.}\end{cases}$
In particular, if $G$ is centerless, then $\operatorname{Good}(\operatorname{Conj}G)=\{\iota\}$.
Proof.
This follows from Corollary 5.3 and Proposition 6.4. ∎
Example 6.6.
Consider the quandle $\operatorname{Conj}X$ from Example 2.16. Although $\operatorname{Conj}X$ is self-dual, Corollary 6.5 states that $\operatorname{Conj}X$ has no good involutions. This can also be seen from the fact that $\operatorname{Conj}X$ is a noninvolutory quandle of order 444; Table 1 shows that no such quandles have good involutions.
6.3. Main results
The following is Theorem 1.3. After proving the theorem, we discuss its interpretation in terms of orbits.
Theorem 6.7.
For all functions $\rho:X\to X$, $\rho$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a function $\zeta:X\to Z(H)$ such that
(6.1) $\rho(x)=(\zeta(x)x)^{-1},\qquad\zeta(yxy^{-1})=\zeta(x)=\zeta\rho(x)$
for all $x,y\in X$.
Proof.
$\implies$” Given $\rho\in\operatorname{Good}(\operatorname{Conj}X)$, Lemma 6.1 states that we can define $\zeta:X\to Z(H)$ by
$\zeta(x):=(\rho(x)x)^{-1}.$
This way, $\rho$ satisfies the first equality in (6.1). Since $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$ (by Proposition 4.2), the second equality in (6.1) follows from Lemma 6.3. For all $x\in X$, the centrality of $\zeta(x)^{-1}=\rho(x)x$ implies that $x\rho(x)=\rho(x)x$, so
$\zeta\rho(x)=(\rho^{2}(x)\rho(x))^{-1}=(x\rho(x))^{-1}=(\rho(x)x)^{-1}=\zeta(x)$
because $\rho$ is an involution. This verifies the final equality in (6.1).
[…]
$\rho^{2}(x)=(\zeta\rho(x)\rho(x))^{-1}=(\zeta(x)\rho(x))^{-1}=(\zeta(x)\zeta(x)^{-1}x^{-1})^{-1}=x$
because $\zeta(x)\in Z(H)$. In other words, $\rho$ is an involution, so by Lemma 6.1, it suffices to show that $\rho\in\operatorname{Anti}(\operatorname{Conj}X)$ and $p(x)x\in Z(H)$ for all $x\in X$. But the former condition follows from Lemma 6.3, and the latter follows from the first equality in (6.1). ∎
Example 6.8.
Taking $\rho$ to be the inversion map $x\mapsto x^{-1}$ corresponds to letting $\zeta:X\to Z(H)$ be the constant function $x\mapsto 1$. Hence, Theorem 6.7 recovers Example 3.13.
Remark 6.9.
The condition that $\zeta(yxy^{-1})=\zeta(x)$ means that $\zeta:X\to Z(H)$ is constant on orbits of $X$ under the action of $H$ by conjugation. These orbits are precisely the connected components of $\operatorname{Conj}X$; cf. Lemma 2.25. Denote the set of all such orbits by $X/H$.
Therefore, Theorem 6.7 states that every good involution $\rho$ of $\operatorname{Conj}X$ factors through the natural projection $\pi:X\twoheadrightarrow X/H$ and a unique function $\zeta^{*}:X/H\to Z(H)$. Moreover, $\zeta^{*}$ completely determines $\rho$.
Remark 6.10.
If $X=G$, then the condition that $\zeta(yxy^{-1})=\zeta(x)$ states that $\zeta:G\to Z(G)$ is a class function.
6.3.1. Corollaries
We provide upper bounds on the number of good involutions of $\operatorname{Conj}X$. In the case that $\operatorname{Conj}X$ is connected, we also simplify the characterization of $\operatorname{Good}(\operatorname{Conj}X)$ in Theorem 6.7. In the following, let $X/H$ denote the set of orbits of $X$ under the action of $H$ by conjugation.
Corollary 6.11.
Suppose that $\operatorname{Conj}X$ is connected. Then for all functions $\rho:X\to X$, $\rho$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a central element $z\in Z(H)$ such that
$\rho(x)=zx^{-1}$
for all $x\in X$. In particular,
$|{\operatorname{Good}(\operatorname{Conj}X)}|\leq|{Z(H)}|.$
If $X$ is not closed under inverses, then this bound is strict.
Proof.
By Remark 6.9, every good involution $\rho$ of $\operatorname{Conj}X$ factors through the natural projection $\pi:X\twoheadrightarrow X/H$ and a unique function $\zeta^{*}:X/H\to Z(H)$. By hypothesis, $X/H=\{X\}$ is a singleton set, so $\zeta^{*}$ is a constant function. Hence, taking
$z:=(\zeta^{*}(X))^{-1}$
yields the main claim due to Theorem 6.7. The final claim follows from Example 6.8. ∎
In other words, if $\operatorname{Conj}X$ is connected, then the set $\operatorname{Good}(\operatorname{Conj}X)$ is in bijection with the set of central elements $z\in Z(H)$ such that $zx^{-1}\in X$ for all $x\in X$.
Even if $\operatorname{Conj}X$ is not connected, the following bound holds.
Corollary 6.12.
Let $k(X):=|{X/H}|$. Then
$|{\operatorname{Good}(\operatorname{Conj}X)}|\leq\min(|{\operatorname{Aut}(\operatorname{Conj}X)}|,|{Z(H)}|^{k(X)}).$
If $X$ is not closed under inverses, then the bound $|{\operatorname{Good}(\operatorname{Conj}X)}|<|{Z(H)}|^{k(X)}$ is strict.
Proof.
The bound $|{\operatorname{Good}(\operatorname{Conj}X)}|\leq|{\operatorname{Aut}(\operatorname{Conj}X)}|$ comes from Corollary 4.8. On the other hand, Remark 6.9 shows that there are at least as many functions $\zeta^{*}:X/H\to Z(H)$ as there are good involutions $\rho$; that is, $|{Z(H)}|^{k(X)}\geq|{\operatorname{Good}(\operatorname{Conj}X)}|$. The final claim follows from Example 6.8. ∎
Remark 6.13.
By Corollary 6.5, the bounds in Corollaries 6.11 and 6.12 are equalities whenever $H$ is centerless and $\operatorname{Conj}X$ is inversion-closed. Furthermore, Example 9.2 provides a connected conjugation subquandle $\operatorname{Conj}X$ of order 303030 that attains these bounds even though $Z(H)$ is nontrivial, and Example 9.5 provides a similar example whose order is infinite.
7. Symmetric core quandles
In this section, we prove Theorem 1.4 (Theorem 7.11) and discuss several related results. This addresses Problems 1.1 and 1.2 for nontrivial core quandles $\operatorname{Core}G$.
Given a group $G$, recall that $\operatorname{Core}G$ is trivial if and only if $G$ is an elementary abelian 222-group. Since good involutions of trivial quandles are already classified (see Example 3.11), we will let $G$ be a group of exponent greater than 222 throughout this section.
7.1. Setup for the proof
To prove Theorem 1.4, we apply Observation 3.9 and Theorem 6.7 to a certain subquandle $\operatorname{Conj}X$ of a conjugation quandle $\operatorname{Conj}K$ such that $\operatorname{Conj}X\cong\operatorname{Core}G$. To study $\operatorname{Conj}X$, we first consider two subgroups of $Z(G)$.
7.1.1. The subgroups $A$ and $T$
Throughout this section, let $A$ be the subset of $Z(G)$ consisting of elements $a\in Z(G)$ satisfying the following condition: There exist a positive integer $n\in\mathbb{Z}^{+}$ and elements $g_{1},\ldots,g_{2n}\in G$ such that
$a=g_{1}g_{2}^{-1}g_{3}g_{4}^{-1}\cdots g_{2n-1}g_{2n}^{-1}=g_{1}^{-1}g_{2}g_{3}^{-1}g_{4}\cdots g_{2n-1}^{-1}g_{2n}.$
Observation 7.1.
$A$ is a subgroup of $Z(G)$ and, hence, a normal subgroup of $G$.
Throughout this section, let $T$ be the subgroup of $Z(G)$ whose elements have order 111 or 222:
$T:=\{z\in Z(G)\mid z^{2}=1\}.$
Proposition 7.2.
$T$ is a subgroup of $A$, and $T=A$ if $G$ is abelian or $Z(G)=T$.
Proof.
Straightforward. ∎
As an aside, we pose the following question.
Problem 7.3.
Does there exist a group $G$ such that $T\neq A$?
7.1.2. Core quandles as conjugation subquandles
Let $K$ be the semidirect product
$K:=(G\times G)\rtimes\mathbb{Z}/2\mathbb{Z},$
where $\mathbb{Z}/2\mathbb{Z}$ acts on $G\times G$ by swapping coordinates. In 2021, Bergman [core]*(6.5) showed that the function $\varphi:G\to K$ defined by
(7.1) $\varphi(g):=(g,g^{-1},1)$
is a quandle embedding of $\operatorname{Core}G$ into $\operatorname{Conj}K$. Henceforth, let $X:=\varphi(G)$ denote the image of $\varphi$, and let $H:=\langle X\rangle$ be the subgroup of $K$ generated by $X$. To prove Theorem 1.4, it suffices to apply Theorem 6.7 to $\operatorname{Conj}X$ by Observation 3.9.
Observation 7.4.
Every element of $X$ has order 222 in $H$.
Observation 7.5.
For all $g\in G$, $H$ contains the element $(g,g^{-1},0)$.
7.1.3. Calculation of the center
To apply Theorem 6.7 to $\operatorname{Conj}X$, we first compute $Z(H)$.
Lemma 7.6.
Let $G$ be a group containing an element $z$ such that $zgz^{-1}=g^{-1}$ for all $g\in G$. Then $G$ is an elementary abelian 222-group.
Proof.
For all $g,h\in G$, conjugating the product $g^{-1}h^{-1}$ by $z$ shows that $g$ and $h$ commute. Thus, $G$ is abelian, so
$g=zgz^{-1}=g^{-1}$
for all $g\in G$. ∎
Lemma 7.7.
Let $a\in Z(G)$. Then $(a,a,0)\in H$ if and only if $a\in A$.
Proof.
[…]
$(a,a,0)=\prod^{2n}_{i=1}(g_{i},g_{i}^{-1},1)=\prod^{2n-1}_{i\text{ odd}}(g_{i}g_{i+1}^{-1},g_{i}^{-1}g_{i+1},0)=(g_{1}g_{2}^{-1}\cdots g_{2n-1}g_{2n}^{-1},g_{1}^{-1}g_{2}\cdots g_{2n-1}^{-1}g_{2n},0)$
in $H$. Since $a\in Z(G)$, the equality of the first two coordinates shows that $a\in A$. ∎
Proposition 7.8.
The center of $H$ is
$Z(H)=\{(a,a,0)\mid a\in A\}\cong A.$
Proof.
If $a\in A$, then $(a,a,0)\in H$ by Lemma 7.7, and verifying that $(a,a,0)\in Z(H)$ is straightforward.
Conversely, let $(a,b,i)\in Z(H)$ be arbitrary. If $i=1$, then Observation 7.5 shows that
$(g,g^{-1},0)=(a,b,1)(g,g^{-1},0)(a,b,1)^{-1}=(ag^{-1}a^{-1},bgb^{-1},0)$
for all $g\in G$. It follows from Lemma 7.6 that $G$ is an elementary abelian 222-group, which contradicts the assumption we made at the beginning of this section.
Thus, we must have $i=0$. By Lemma 7.7, it now suffices to show that $a\in Z(G)$ and $a=b$. Indeed,
$(g,g^{-1},0)=(a,b,0)(g,g^{-1},0)(a,b,0)^{-1}=(aga^{-1},bg^{-1}b^{-1},0)$
for all $g\in G$, so $a\in Z(G)$. On the other hand,
$(1,1,1)=(a,b,0)(1,1,1)(a,b,0)^{-1}=(ab^{-1},ba^{-1},1),$
so $a=b$. ∎
7.1.4. Preliminary versions of Theorem 1.4
We state two preliminary versions of Theorem 1.4. Denote the rack structure of $\operatorname{Conj}X$ by $t$.
Proposition 7.9.
For all functions $\rho^{\prime}:X\to X$, $\rho^{\prime}$ is a good involution of $\operatorname{Conj}X$ if and only if there exists a function $\zeta:X\to Z(H)$ such that
$\rho^{\prime}(g,g^{-1},1)=\zeta(g,g^{-1},1)(g,g^{-1},1),\qquad\zeta t_{(g,g^{-1},1)}=\zeta=\zeta\rho^{\prime}$
for all $(g,g^{-1},1)\in X$.
Proof.
This follows from combining Theorem 6.7 with Observation 7.4. ∎
In the following, denote the rack structure of $\operatorname{Core}G$ by $s$.
Proposition 7.10.
For all functions $\rho:G\to G$, $\rho$ is a good involution of $\operatorname{Core}G$ if and only if there exists a function $\alpha:G\to A$ such that
$\rho(g)=\alpha(g)g,\qquad\alpha s_{g}=\alpha=\alpha\rho$
for all $g\in G$.
Proof.
We will show that the claim is equivalent to Proposition 7.9. Given any function $\alpha:G\to A$, define a function $\mathcal{F}(\alpha):X\to H$ by
$\mathcal{F}(\alpha)(g,g^{-1},1):=(\alpha(g),\alpha(g),0).$
By Proposition 7.8, the image of $\mathcal{F}(\alpha)$ lies in $Z(H)$. Conversely, given any function $\zeta:X\to Z(H)$, define a function $\mathcal{F}^{-1}(\zeta):G\to G$ by sending $g\in G$ to the first coordinate of $\zeta(g,g^{-1},1)\in Z(H)$. By Proposition 7.8, the image of $\mathcal{F}^{-1}(\zeta)$ lies in $A$. Hence, it is straightforward to verify that the assignments $\mathcal{F}$ and $\mathcal{F}^{-1}$ are mutually inverse.
Now, recall that the map $\varphi$ from (7.1) is a quandle isomorphism from $\operatorname{Core}G$ to $\operatorname{Conj}X$. The equivalence of the claim with Proposition 7.9 therefore follows from Observation 3.9 and the following straightforward calculations: For all elements $g\in G$ and functions $\rho:G\to G$ and $\rho^{\prime}:X\to X$ respectively satisfying the conditions of the claim and Proposition 7.9, we have
$\varphi\rho\varphi^{-1}(g,g^{-1},1)=\mathcal{F}(\alpha)(g,g^{-1},1)(g,g^{-1},1),\qquad\varphi^{-1}\rho^{\prime}\varphi(g)=\mathcal{F}^{-1}(\zeta)(g)g,$
so that
$\mathcal{F}(\alpha s_{g})=\mathcal{F}(\alpha)t_{(g,g^{-1},1)},\qquad\mathcal{F}(\alpha\rho)=\mathcal{F}(\alpha)\varphi\rho\varphi^{-1}$
and, similarly,
$\mathcal{F}^{-1}(\zeta t_{(g,g^{-1},1)})=\mathcal{F}^{-1}(\zeta)s_{g},\qquad\mathcal{F}^{-1}(\zeta\rho^{\prime})=\mathcal{F}^{-1}(\zeta)\varphi^{-1}\rho^{\prime}\varphi.$
7.2. Main results
We now prove Theorem 1.4. Preserve the notation from the previous subsection, so that $T$ denotes the subgroup $T=\{z\in Z(G)\mid z^{2}=1\}$.
Theorem 7.11.
Proposition 7.10 holds when $A$ is replaced with $T$.
Proof.
If $\alpha:G\to T$ is a function satisfying the conditions of Proposition 7.10, then $\alpha$ induces a good involution $\rho$ by Propositions 7.2 and 7.10.
Conversely, let $\rho$ be a good involution of $\operatorname{Core}G$. Then $\rho$ is induced by a unique function ${\alpha:G\to A}$ satisfying the conditions of Proposition 7.10, and for all $g\in G$, we have
$s_{\rho(g)}=s_{g}^{-1}=s_{g}$
because $\operatorname{Core}G$ is involutory. Hence, [core2]*Lem. 4.1 states that
$T\ni g^{-1}\rho(g)=g^{-1}\alpha(g)g=\alpha(g),$
where we have used the fact that $\alpha(g)\in Z(G)$. Since $g\in G$ was arbitrary, the image of $\alpha$ lies in $T$, as desired. ∎
Example 7.12.
The constant functions $\alpha:G\to T$ recover Example 3.14.
Remark 7.13.
Let $\mathcal{O}:=G/\operatorname{Inn}(\operatorname{Core}G)$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$. Then Theorem 7.11 states that every good involution $\rho$ of $\operatorname{Core}G$ factors through the natural projection $\pi:G\twoheadrightarrow\mathcal{O}$ and a unique function $\alpha^{*}:\mathcal{O}\to T$, and $\alpha^{*}$ completely determines $\rho$.
7.2.1. Corollaries
Using Remark 7.13, we obtain the following special cases of Theorem 7.11.
Corollary 7.14.
Let $\mathcal{O}:=G/\operatorname{Inn}(\operatorname{Core}G)$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$. Then the following inequality holds:
$|T|\leq|{\operatorname{Good}(\operatorname{Core}G)}|\leq|T|^{|\mathcal{O}|}.$
Proof.
The lower bound in Corollary 7.14 is obtained from the constant functions $\alpha^{*}:\mathcal{O}\to T$; cf. Example 7.12. The upper bound is a consequence of Remark 7.13. ∎
Remark 7.15.
In Propositions 8.68.8, we show that infinitely many groups $G$ attain the upper bound in Corollary 7.14 while exceeding the lower bound. On the other hand, the next two corollaries consider cases in which the lower and upper bounds are equal.
Corollary 7.16.
Suppose that $G$ is generated by the squares of its elements; that is, $G=\langle g^{2}\mid g\in G\rangle$. Then good involutions $\rho$ of $\operatorname{Core}G$ are precisely the multiplication maps $g\mapsto tg$ with $t\in T$. In particular, $\operatorname{Good}(\operatorname{Core}G)$ is in bijection with $T$.
Proof.
By Lemma 2.26, the hypothesis is equivalent to the statement that $\operatorname{Core}G$ is connected. That is, the orbit decomposition $\mathcal{O}$ in Corollary 7.14 is a singleton set, so the claim follows. ∎
Corollary 7.17.
If $Z(G)$ is 222-torsionless, then the only good involution of $\operatorname{Core}G$ is the identity map $\operatorname{id}_{G}$.
Proof.
This follows immediately from Theorem 7.11. ∎
Remark 7.18.
Corollary 7.17 can also be obtained from Example 2.7 and Corollary 5.5.
Remark 7.19.
If $G$ is abelian, then Theorem 7.11 and Corollary 7.16 specialize to characterizations of good involutions of Takasaki kei $T(G)$. In particular, Corollary 7.16 applies to all Takasaki kei of 222-divisible abelian groups; cf. [flat]. Moreover, Corollary 7.17 vastly generalizes a result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles of odd order; cf. Example 5.6.
8. Algorithms
As applications of Theorems 6.7 and 7.11, we describe procedures for computing $\operatorname{Good}R$ when $R$ is a conjugation subquandle $\operatorname{Conj}X$ or a core quandle $\operatorname{Core}G$. We also discuss computational findings for groups up to order 232323.
Carrying over our notation from Section 6, let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle\leq G$.
8.1. How to compute $\operatorname{Good}(\operatorname{Conj}X)$
Iterating the following procedure constructs all good involutions $\rho$ of $\operatorname{Conj}X$. The procedure is based on Theorem 6.7; cf. Remark 6.9.
Algorithm 8.1.
Let $X/H$ be the set of orbits of $X$ under the action of $H$ by conjugation (that is, the canonical action of $\operatorname{Inn}X$), and let $\pi:X\twoheadrightarrow X/H$ be the natural projection.
[…]
    First, pick any function $\zeta^{*}:X/H\to Z(H)$, and define functions $\zeta:X\to Z(H)$ and $\rho:X\to H$ by
    $\zeta:=\zeta^{*}\pi,\qquad\rho(x):=(\zeta(x)x)^{-1}.$
[…]
    If $\rho(X)\subseteq X$ and $\zeta\rho=\zeta$, then Theorem 1.3 states that $\rho$ is a good involution, so return $\rho$. Else, return to the previous step and pick a different function $\zeta^{*}$.
To compute $\operatorname{Good}(\operatorname{Conj}X)$, repeat these steps until all possible functions $\zeta^{*}$ have been considered.
Theorem 6.7 states that $\operatorname{Good}(\operatorname{Conj}X)$ is precisely the set of involutions $\rho$ that can be produced by the above method; cf. Remark 6.9.
Remark 8.2.
If $\operatorname{Conj}X$ is connected, then Algorithm 8.1 simplifies via Corollary 6.11 in the obvious way: For each central element $z\in Z(H)$, test whether $zx^{-1}\in X$ for all $x\in X$. The good involutions $\rho$ of $\operatorname{Conj}X$ are precisely the maps $x\mapsto zx^{-1}$ for which this test succeeds.
8.1.1. Example of Algorithm 8.1
In the following example, we use Algorithm 8.1 to produce a good involutions of $\rho$ of a conjugation quandle such that $\rho$ is neither the inversion map nor the identity map.
Example 8.3.
[…]
$D_{n}=\langle r,s\mid r^{n}=1=s^{2},\ srs=r^{-1}\rangle.$
Following Algorithm 8.1, we construct a good involution $\rho$ of $\operatorname{Conj}D_{n}$.
Since $n$ is even, the center of $D_{n}$ is $Z(D_{n})=\{1,r^{n/2}\}$. Define the function $\zeta^{*}$ on conjugacy classes $C$ of $D_{n}$ by
$\zeta^{*}(C):=\begin{cases}1&\text{if }C\subset Z(D_{n}),\\ r^{n/2}&\text{otherwise,}\end{cases}$
so that $\zeta:=\zeta^{*}\pi$ is defined on elements $g$ of $D_{n}$ by
$\zeta(g)=\begin{cases}1&\text{if }g\in Z(D_{n}),\\ r^{n/2}&\text{otherwise.}\end{cases}$
Then $\rho:D_{n}\to D_{n}$ is the function defined by
$\rho(g):=(\zeta(g)g^{-1})^{-1}=\begin{cases}g&\text{if }g\in Z(D_{n}),\\ r^{n/2-k}&\text{if }g=r^{k}\text{ where }1\leq k<n\text{ and }k\neq\frac{n}{2},\\ sr^{n/2+k}&\text{if }g=sr^{k}\text{ for some }0\leq k<n.\end{cases}$
Verifying that $\zeta\rho=\zeta$ is straightforward. Hence, Theorem 6.7 states that $\rho$ is a good involution of $\operatorname{Conj}D_{n}$. Note that $\rho$ does not equal the inversion map $g\mapsto g^{-1}$.
Combined with Corollary 6.5 (or Corollary 5.3), Example 8.3 yields the following.
Proposition 8.4.
Let $n\geq 3$ be an integer. Then the conjugation quandle $\operatorname{Conj}D_{n}$ has more than one good involution if and only if $n$ is even.
8.2. How to compute $\operatorname{Good}(\operatorname{Core}G)$
Let $G$ be a group of exponent greater than 222, so that $\operatorname{Core}G$ is a nontrivial quandle. Similarly to the procedure for $\operatorname{Conj}X$, the following procedure constructs all good involutions $\rho$ of $\operatorname{Core}G$. The procedure is based on Theorem 7.11; cf. Remark 7.13.
Algorithm 8.5.
Let $\mathcal{O}$ be the set of orbits of $G$ under the action of $\operatorname{Inn}(\operatorname{Core}G)$, and let $\pi:G\twoheadrightarrow\mathcal{O}$ be the natural projection. Also, let $T:=\{z\in Z(G)\mid z^{2}=1\}$.
[…]
    First, pick any function $\alpha^{*}:\mathcal{O}\to T$, and define functions $\alpha:G\to T$ and $\rho:G\to G$ by
    $\alpha:=\alpha^{*}\pi,\qquad\rho(g):=\alpha(g)g.$
[…]
    If $\alpha\rho=\alpha$, then Theorem 1.4 states that $\rho$ is a good involution of $\operatorname{Core}G$, so return $\rho$. Else, return to the previous step and pick a different function $\alpha^{*}$.
To compute $\operatorname{Good}(\operatorname{Core}G)$, repeat these steps until all possible functions $\alpha^{*}$ have been considered.
Theorem 7.11 states that $\operatorname{Good}(\operatorname{Core}G)$ is precisely the set of involutions $\rho$ that can be produced by the above method; cf. Remark 7.13.
8.2.1. Examples of Algorithm 8.5
The following three examples show that the upper bound in Corollary 7.14 is sharp for infinitely many nontrivial, nonconnected core quandles.
[…]
$\operatorname{Dic}_{n}=\langle a,x\mid a^{2n}=1,\ x^{2}=a^{n},\ xax^{-1}=a^{-1}\rangle.$
Proposition 8.6.
For all odd integers $n\geq 3$, the core quandle $\operatorname{Core}\operatorname{Dic}_{n}$ has exactly $2^{(n+3)/2}$ good involutions. In particular, $\operatorname{Dic}_{n}$ exceeds the lower bound and attains the upper bound in Corollary 7.14.
Proof.
[…]
  • one orbit $\pi(x)=\{x,x^{3}\}$ of size 222,
[…]
  • the following $(n-1)/2$ orbits of size 444:
    $\pi(xa^{m})=\{xa^{m},xa^{m+1},xa^{2n-m-1},xa^{2n-m}\}\qquad\text{for all }1\leq m\leq\frac{n-1}{2}.$
In particular, the lower and upper bounds in Corollary 7.14 equal 222 and $2^{(n+3)/2}$, respectively.
Evidently, each orbit in $\mathcal{O}$ is closed under multiplication by elements of $T$. That is,
$\pi(tg)=\pi(g)$
for all $t\in T$ and $g\in\operatorname{Dic}_{n}$. In particular, given any function $\alpha^{*}:\mathcal{O}\to T$, letting $g\in G$ be arbitrary and taking $t:=\alpha^{*}\pi(g)$ yields
$\pi(\alpha^{*}\pi(g)g)=\pi(g)$
for all $g\in G$. Applying $\alpha^{*}$ to both sides recovers the condition that $\alpha\rho=\alpha$ in Algorithm 8.5. Hence, Theorem 7.11 states that $\alpha^{*}$ defines a unique good involution $\rho$ of $\operatorname{Core}\operatorname{Dic}_{n}$.
Since $\alpha^{*}:\mathcal{O}\to T$ was an arbitrary function, Theorem 7.11 states that the $2^{(n+3)/2}$ possible functions $\alpha^{*}$ collectively define $2^{(n+3)/2}$ distinct good involutions $\rho$ of $\operatorname{Core}D_{n}$, and no other good involutions exist. This completes the proof. ∎
Similar arguments show the following.
Proposition 8.7.
For all even integers $n\geq 2$, the core quandle $\operatorname{Core}\operatorname{Dic}_{n}$ has exactly $2^{n/2+3}$ good involutions. In particular, $\operatorname{Dic}_{n}$ exceeds the lower bound and attains the upper bound in Corollary 7.14.
Proposition 8.8 (cf. [symm-quandles-2]*Thm. 3.2).
If $n\geq 4$ is a multiple of 444, then the dihedral quandle $R_{n}=T(\mathbb{Z}/n\mathbb{Z})$ has exactly four good involutions. In particular, $\mathbb{Z}/n\mathbb{Z}$ exceeds the lower bound and attains the upper bound in Corollary 7.14.
8.3. Enumeration
We implemented Algorithms 8.1 and 8.5 in GAP [GAP4] and used them to compute $\operatorname{Good}(\operatorname{Conj}G)$ and $\operatorname{Good}(\operatorname{Core}G)$ for various groups $G$ up to order 232323. We provide our code and raw data (including the explicit mappings $\rho:G\to G$) in a GitHub repository [code].
Table 2 counts good involutions of conjugation quandles $\operatorname{Conj}G$ of nonabelian groups up to order 222222. Table 3 does the same for core quandles $\operatorname{Core}G$ of groups of exponent greater than 222 up to order 232323. In Appendix B, we specify these counts for each group $G$ considered in Tables 2 and 3.
Table 2. Number of good involutions across all nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order 222222. Cf. Table 6.
Order of $G$ 6 8 10 12 14 16 18 20 21 22
Good involutions 1 32 1 17 1 13056 66 33 1 1
Table 3. Number of good involutions across all nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323. Cf. Table LABEL:tab:app2.
Order of $G$ 3 4 5 6 7 8 9 10 11 12 13
Good involutions 1 4 1 3 1 72 2 3 1 31 1
Order of $G$ 14 15 16 17 18 19 20 21 22 23
Good involutions 3 1 10856 1 7 1 47 2 3 1
9. Constructions
In this section, we compute $\operatorname{Good}(\operatorname{Conj}X)$ for various connected conjugation subquandles $\operatorname{Conj}X$. This addresses Problem 1.2. By considering examples in which $\operatorname{Conj}X$ is inversion-closed, we also construct infinite families of CNS-quandles. This addresses Problems 1.6 and 1.7.
Carrying over our notation from Section 6, let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle\leq G$. Also, let $\iota:G\to G$ denote the inversion map $g\mapsto g^{-1}$.
9.1. Constructing CNS-quandles
The following provides a robust, group-theoretic way to construct CNS-quandles.
Proposition 9.1.
Suppose that $\operatorname{Conj}X$ satisfies the following three conditions:
  • $X$ is a conjugacy class of $H$.
  • There exists an element $x\in X$ such that $x^{2}\notin Z(H)$.
  • $\operatorname{Conj}X$ is inversion-closed.
Then $(\operatorname{Conj}X,\iota|_{X})$ is a CNS-quandle.
Proof.
This follows from Lemma 2.25, Observation 2.11, and Example 3.13. ∎
9.2. Examples
In the following examples, we study good involutions of connected subquandles of special linear groups $\mathrm{SL}(m,q)$, symmetric groups $S_{n}$, and the special unitary group $\mathrm{SU}(2)$.
Example 9.2.
Let $G$ be the special linear group $\mathrm{SL}(2,5)$. We study good involutions of several connected subquandles of $\operatorname{Conj}G$. Using GAP [GAP4], we find that $G$ has seven conjugacy classes $X$ of size greater than 111: four of size 121212, two of size 202020, and one of size 303030.
It is well-known that the only nontrivial normal subgroup of $G$ is the center
$Z(G)=\{\pm I_{2}\}.$
It follows that each of the seven aforementioned conjugacy classes $X$ generates $G$, so by Lemma 2.25, all of them are connected subquandles $\operatorname{Conj}X$ of $\operatorname{Conj}G$. Using GAP, we verify that the subquandles of orders 121212 and 202020 satisfy the conditions in Proposition 9.1. Therefore, $\operatorname{Conj}G$ contains four CNS-subquandles $(\operatorname{Conj}X,\iota|_{X})$ of order 121212 and two CNS-subquandles of order 202020.
On the other hand, the subquandle $\operatorname{Conj}X$ of order 303030 satisfies only the first two conditions in Proposition 9.1, so $(\operatorname{Conj}X,\iota|_{X})$ is a connected symmetric kei. Since the element $-I_{2}\in Z(G)$ corresponds to the identity map in Corollary 6.11, it follows from Proposition 3.16 and Corollary 6.11 that
$\operatorname{Good}(\operatorname{Conj}X)=\begin{cases}\{\iota|_{X},\operatorname{id}_{X}\}&\text{if }|X|=30,\\ \{\iota|_{X}\}&\text{if }|X|=12,20.\end{cases}$
Example 9.3.
Let $m\geq 3$ be an integer, and let $\mathbb{F}_{q}$ be a finite field of order $q$ and characteristic $p\geq 3$. Let $G$ be the special linear group $\mathrm{SL}(m,q)$, and let $X\subset G$ be the conjugacy class of the matrix
$M:=\begin{pmatrix}1&1&\\ &1&\\ &&I_{n-2}\end{pmatrix}.$
Since $m\geq 3$, $X$ is the set of all transvections (that is, nonidentity linear transformations of $\mathbb{F}_{q}^{m}$ that fix a hyperplane) in $G$, which is a symmetric generating subset of $G$; see, for example, [suzuki]*Ch. 1, Sec. 9. In particular,
$Z(H)=Z(G)=\{\lambda I_{n}\mid\lambda^{n}=1\},$
so we have
$M^{2}=\begin{pmatrix}1&2&\\ &1&\\ &&I_{n-2}\end{pmatrix}\notin Z(H)$
[…]
Finally, we show that $\iota|_{X}$ is the only good involution of $\operatorname{Conj}X$. By Corollary 6.11 and Example 6.8, it suffices to show that there exists a transvection $B\in X$ such that for all transvections $A\in X$, the product $AB$ does not equal any nonidentity element $\lambda I$ of $Z(H)$. But this follows from the fact that $A$ and $B$ are unipotent (due to being transvections).
Example 9.4.
Let $n\geq 5$ be an integer, and let $X$ be any conjugacy class in the symmetric group $S_{n}$ satisfying one of the following conditions:
  • $X$ contains odd permutations; that is, $X\cap A_{n}=\emptyset$.
  • $X$ is a single conjugacy class in $A_{n}$; that is, $X$ does not split into two conjugacy classes in the alternating group $A_{n}$.
If $X$ does not contain involutions, then a straightforward verification of the conditions in Proposition 9.1 shows that $(\operatorname{Conj}X,\iota|_{X})$ is a CNS-quandle. In fact, since $S_{n}$ and $A_{n}$ are centerless, Corollary 6.5 shows that $\iota|_{X}$ is the only good involution of $\operatorname{Conj}X$.
In particular, let $k$ be an integer such that $3\leq k\leq 2\lfloor n/2\rfloor$, and let $X$ be the set of $k$-cycles in $S_{n}$. Then we obtain the second part of Proposition 1.8.
Table 4. Unique sizes of conjugacy classes $X$ of $S_{n}$ that satisfy the conditions in Example 9.4.
$n$ Unique sizes of $X$
555 202020, 303030
666 404040, 909090, 120120120
777 707070, 210210210, 280280280, 420420420, 504504504, 630630630, 840840840
888 112112112, 420420420, 112011201120, 126012601260, 134413441344, 168016801680, 252025202520, 336033603360, 403240324032, 504050405040
999
168168168, 756756756, 224022402240, 252025202520, 302430243024, 336033603360, 756075607560, 907290729072, 100801008010080,
113401134011340, 151201512015120, 181441814418144, 201602016020160, 259202592025920, 302403024030240, 453604536045360
Addressing Problem 1.7, Table 4 displays the orders of the CNS-subquandles $(\operatorname{Conj}X,\iota|_{X})$ of $(\operatorname{Conj}S_{n},\iota)$ constructed in Example 9.4 for all $5\leq n\leq 9$, omitting duplicate orders. These orders were obtained using the GAP [GAP4] program provided in Appendix A.
9.2.1. An infinite-order example
While the previous examples have studied finite conjugation subquandles, our methods also apply in the infinite-order case.
Example 9.5.
[…]
It is well-known (see, for example, [hall]*Ex. 9 in Ch. 11, Sec. 8) that the diagonal matrices
$d_{\theta}:=\operatorname{diag}(e^{i\theta},e^{-i\theta})$
with $\theta\in[0,\pi]$ represent all conjugacy classes of $G$, and an element $g\in G$ is conjugate to $d_{\theta}$ if and only if the eigenvalues of $g$ are $e^{\pm i\theta}$. Let $X_{\theta}$ denote the conjugacy class of $d_{\theta}$ in $G$.
Fix $\theta\in(0,\pi)$. Since the only nontrivial normal subgroup of $G$ is the center
$Z(G)=\{\pm I_{2}\},$
it follows that $X_{\theta}$ generates $G$. Thus, Lemma 2.25 shows that $\operatorname{Conj}X_{\theta}$ is connected. By the above discussion, $\operatorname{Conj}X_{\theta}$ is also inversion-closed. By verifying the conditions in Corollary 6.11 and Proposition 9.1, we deduce that
$\operatorname{Good}(\operatorname{Conj}X_{\theta})=\begin{cases}\{\iota|_{X_{\theta}},\operatorname{id}_{X_{\theta}}\}&\text{if }\theta=\pi/2,\\ \{\iota|_{X_{\theta}}\}&\text{otherwise,}\end{cases}$
and $(\operatorname{Conj}(X_{\theta}),\iota|_{X_{\theta}})$ is a CNS-quandle if and only if $\theta\neq\pi/2$. In particular, $(\operatorname{Conj}G,\iota)$ contains uncountably many CNS-subquandles.
10. Connections with Legendrian racks
In this section, we prove Theorem 1.10 and several other equivalences of categories (Theorem 10.1). Before that, we introduce kink-involutory racks and discuss medial racks, GL-racks, and Legendrian racks. We also discuss a canonical rack automorphism $\theta$ that plays a fundamental role in the theory.
10.1. Discussion of results
The following is the main result of this section.
Theorem 10.1.
In each of the following pairs or triples, the categories mentioned are equivalent (in fact, isomorphic), and so are their full subcategories whose objects are medial:
  • Racks $(\mathsf{Rack})$ and Legendrian racks $(\mathsf{LR})$.
  • Racks for which the canonical automorphism $\theta$ is an involution $(\mathsf{KIR})$ and Legendrian quandles $(\mathsf{LQ})$.
  • Involutory GL-racks $(R,\mathtt{u})$ in which $\mathtt{u}$ is an involution and involutory symmetric racks.
  • Involutory racks $(\mathsf{InvRack})$, Legendrian kei $(\mathsf{LK})$, and symmetric kei $(\mathsf{SK})$.
After we construct functors between $\mathsf{Rack}$ and $\mathsf{LR}$, Theorem 10.1 will more or less follow from the definitions, all of which we provide in this section.
10.1.1. Corollaries
In tandem with [taGL]*Thm. 5.6, Theorem 10.1 yields the following.
Corollary 10.2.
The category $\mathsf{LR}$ and the category of GL-quandles are isomorphic, and so are their full subcategories whose objects are medial.
By combining Theorem 10.1 with [taGL]*Thm. 5.1, we deduce the center $Z(\mathsf{SK})$ of the category of symmetric kei.
Corollary 10.3.
The center of the category $\mathsf{SK}$ is the group generated by the collection of all good involutions:
$Z(\mathsf{SK})=\langle\rho\mid\rho^{2}=1\rangle\cong\mathbb{Z}/2\mathbb{Z}.$
Since $\mathsf{LK}$ and $\mathsf{SK}$ are isomorphic, classification results in either category hold in the other. For example, Theorem 7.11 classifies all Legendrian structures on core quandles $\operatorname{Core}G$, and (by Observation 2.11) Theorem 6.7 classifies all Legendrian structures on subquandles $\operatorname{Conj}X$ of conjugation quandles $\operatorname{Conj}G$ such that $x^{2}\in Z(\langle X\rangle)$ for all $x\in X$.
Conversely, [taGL]*Prop. 4.11 slightly refines a classification result of Kamada and Oshiro [symm-quandles-2]*Thm. 3.2 for dihedral quandles of even order. Note that Corollary 7.17 recovers their result for dihedral quandles of odd order.
Corollary 10.4.
[…]
$\operatorname{Good}R_{2n}=\begin{cases}\{\operatorname{id}_{\mathbb{Z}/2n\mathbb{Z}},\alpha_{1,n}\}&\text{if }n\text{ is odd,}\\ \{\operatorname{id}_{\mathbb{Z}/2n\mathbb{Z}},\alpha_{1,n},\alpha_{n+1,0},\alpha_{n+1,n}\}&\text{otherwise.}\end{cases}$
Of the corresponding symmetric kei, the only isomorphic ones are
$(R_{2n},\alpha_{n+1,0})\cong(R_{2n},\alpha_{n+1,n})$
for even $n$. In particular, there are two isomorphism classes of symmetric kei of the form $(R_{2n},\rho)$ when $n$ is odd, and there are three when $n$ is even.
10.1.2. Computer search
Table 5. Enumeration of isomorphism classes of various types of Legendrian racks up to order 888.
Order 0 1 2 3 4 5 6 7 8
Legendrian racks 1 1 2 6 19 74 353 2080 16023
Medial Legendrian racks 1 1 2 6 18 68 329 1965 15455
Legendrian quandles 1 1 2 5 15 54 240 1306 9477
Medial Legendrian quandles 1 1 2 5 14 48 219 1207 9042
Legendrian kei 1 1 2 5 13 42 180 906 6317
Medial Legendrian kei 1 1 2 5 12 38 168 850 6090
Adding to our classification of symmetric racks up to order 888 using GAP [GAP4], we used the classification of GL-racks up to order 888 in [taGL] to classify various types of Legendrian racks up to isomorphism.
We provide our code and data in a GitHub repository [gl-code]. Table 5 enumerates our data; cf. Table 1, [library]*Tab. 1, and [taGL]*Tab. A.1. Note that the rows of Table 5 also count isomorphism classes of the other objects listed in Theorem 10.1, some of which were recorded in [library]*Tab. 1.
10.2. Preliminaries: Racks
[…]
10.2.1. The canonical automorphism $\theta$
Every rack $R=(X,s)$ has a canonical automorphism $\theta_{R}$ defined by
$x\mapsto s_{x}(x)$
for all $x\in X$; see [center]*Prop. 2.5 and cf. [taGL]*Prop. 2.16.333Some authors, especially those working in framed knot theory (e.g., [quandlebook]*p. 149), call $\theta_{R}$ the kink map of $R$ and denote it by $\pi$. This motivates our choice of nomenclature in Definition 10.7. Evidently, $R$ is a quandle if and only if $\theta_{R}=\operatorname{id}_{X}$, so we can loosely think of $\theta_{R}$ as the obstruction for $R$ to be a quandle. When there is no ambiguity, we will suppress the subscript and only write $\theta$ to mean $\theta_{R}$.
Example 10.5.
If $(X,s)$ is a permutation rack defined by $\sigma\in S_{X}$, then $\theta=\sigma$.
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. Concretely, $\eta\in Z(\mathcal{C})$ if and only if, for all objects $R,S$ and morphisms $\varphi:R\to S$ in $\mathcal{C}$, the component $\eta_{R}$ is an endomorphism of $R$, and
$\eta_{S}\varphi=\varphi\eta_{R}.$
Our proof of Theorem 10.1 uses the following result of Szymik [center] in 2018.
Proposition 10.6 ([center]*Thm. 5.4).
The center $Z(\mathsf{Rack})$ of the category of racks is the free group $\langle\theta\rangle\cong\mathbb{Z}$ generated by the collection $\theta$ of canonical automorphisms $\theta_{R}$ for all racks $R$.
10.2.2. Kink-involutory racks
We introduce a class of racks called kink-involutory racks that generalize quandles and involutory racks.
Definition 10.7.
A rack $R=(X,s)$ is called kink-involutory if its canonical automorphism $\theta$ is an involution. Let $\mathsf{KIR}$ be the full subcategory of $\mathsf{Rack}$ whose objects are kink-involutory.
Example 10.8.
All quandles are kink-involutory.
Proposition 10.9.
All involutory racks are kink-involutory.
Proof.
Let $R=(X,s)$ be an involutory rack. For all $x\in X$,
$\theta^{2}(x)=\theta s_{x}(x)=s_{s_{x}(x)}s_{x}(x)=s_{x}s_{x}(x)=x.$
Thus, $\theta^{2}=\operatorname{id}_{X}$. ∎
Example 10.10.
For an example of a kink-involutory rack that is neither involutory nor a quandle, consider the rack $R=(X,s)$ with underlying set $X=\{1,2,3,4,5\}$ whose rack structure is given by the permutations
$s_{1}=s_{2}=(12)(345),\quad\quad s_{3}=s_{4}=s_{5}=(12)$
in cycle notation. Then $s_{1}^{2}\neq\operatorname{id}_{X}$, so $R$ is not involutory. Also, $\theta=(12)$ is a nonidentity involution, so $R$ is kink-involutory and not a quandle.
10.2.3. Medial racks
We discuss a class of racks called medial racks or abelian racks, which Joyce [joyce] introduced to study Alexander invariants of knots. [^4] In this paper, we adopt the name “medial” over “abelian.” This is to prevent confusion with commutative racks, which satisfy the much rarer condition that $s_{x}(y)=s_{y}(x)$ for all $x,y\in X$. Certain categorical and universal-algebraic notions (see, for example, [rack-roll]*Sec. 3, [Hom]*Thm. 12, or [taGL]*Sec. 2.3.3) allow medial racks to enhance coloring invariants of smooth knots [Hom]*Ex. 9 and Legendrian knots [taGL]*Cor. 3.21.
Definition 10.11 ([rack-roll]*Sec. 3).
Let $R=(X,s)$ be a rack, and consider the product rack $R\times R$. We say that $R$ is medial or abelian if the map $X\times X\to X$ defined by
$(x,y)\mapsto s_{y}(x)$
is a rack homomorphism from $R\times R$ to $R$.
Remark 10.12.
Unpacking Definition 10.11 recovers the pointwise definition of mediality introduced by Joyce [joyce]*Def. 1.3.
Example 10.13.
All permutation racks and Takasaki kei are medial.
10.3. Preliminaries: GL-racks
We discuss GL-racks, which Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced in 2023 to develop invariants of Legendrian knots in contact three-space. See [taGL] for a development of the algebraic theory of GL-racks and [he] for applications.
While the following definition contrasts with the original definitions of Karmakar et al. and Kimura, their equivalence was proven in [taGL]*Prop. 3.12.
Definition 10.14 ([taGL]).
Given a rack $R=(X,s)$, a GL-structure on $R$ is a rack automorphism $\mathtt{u}\in\operatorname{Aut}R$ such that $\mathtt{u}s_{x}=s_{x}\mathtt{u}$ for all $x\in X$. We call the pair $(R,\mathtt{u})$ a GL-rack, generalized Legendrian rack, or bi-Legendrian rack.
Example 10.15 ([bi]*Ex. 3.7).
Given a permutation rack $P=(X,s)$ defined by $\sigma\in S_{X}$, the set of GL-structures on $P$ is precisely the centralizer $C_{S_{X}}(\sigma)$.
Definition 10.16.
A GL-rack homomorphism between two GL-racks $(R_{1},\mathtt{u}_{1})$ and $(R_{2},\mathtt{u}_{2})$ is a rack homomorphism $\varphi$ from $R_{1}$ to $R_{2}$ that satisfies $\varphi\mathtt{u}_{1}=\mathtt{u}_{2}\varphi$. Let $\mathsf{GLR}$ be the category of GL-racks.
10.3.1. Centrality of GL-structures
By Definition 10.16, all integer powers of GL-structures $\mathtt{u}$ and canonical rack automorphisms $\theta$ lie in the categorical center $Z(\mathsf{GLR})$. In fact, $Z(\mathsf{GLR})\cong\mathbb{Z}^{2}$ is the free abelian group generated by these two collections of automorphisms; see [taGL]*Thm. 5.1.
10.3.2. Legendrian racks
As their name suggests, Legendrian racks are a special class of GL-racks defined below. They were introduced by Ceniceros et al. [ceniceros] in 2021, and they generalize the constructions of Kulkarni and Prathamesh [original].
Although the following definition differs from the original definition of Ceniceros et al., their equivalence was proven in [taGL]*Cor. 3.13.
Definition 10.17 ([taGL]).
Let $(R,\mathtt{u})$ be a GL-rack. We say that $(R,\mathtt{u})$ is a Legendrian rack if $\theta=\mathtt{u}^{-2}$, in which case we say that $\mathtt{u}$ is a Legendrian structure on $R$.
We say that a Legendrian rack $(R,\mathtt{u})$ is a Legendrian quandle if $R$ is a quandle or, equivalently, if $\mathtt{u}$ is an involution. Similarly, we say that $(R,\mathtt{u})$ is a Legendrian kei if $R$ is a kei.
Let $\mathsf{LR}$, $\mathsf{LQ}$, and $\mathsf{LK}$ be the full subcategories of $\mathsf{GLR}$ whose objects are Legendrian racks, Legendrian quandles, and Legendrian kei, respectively.
Example 10.18 ([bi]*Ex. 3.6).
Let $G$ be a group, and let $z\in Z(G)$ be a central element of $G$. Then multiplication by $z$ defines a GL-structure on the conjugation quandle $\operatorname{Conj}G$. This GL-quandle is a Legendrian quandle if and only if $z^{2}=1$ in $G$.
The following part of Theorem 10.1 follows immediately.
Proposition 10.19.
In each of the following pairs, the categories mentioned have precisely the same objects and morphisms:
  • Involutory GL-racks $(R,\mathtt{u})$ in which $\mathtt{u}$ is an involution and involutory symmetric racks.
  • Legendrian kei $(\mathsf{LK})$ and symmetric kei $(\mathsf{SK})$.
Proof.
This follows directly from Proposition 3.18 and the definitions of $\mathsf{GLR}$ and $\mathsf{LK}$. ∎
10.4. Proof of Theorem 10.1
To prove Theorem 10.1, we construct functors $F:\mathsf{LR}\to\mathsf{Rack}$ and $F^{-1}:\mathsf{Rack}\to\mathsf{LR}$ satisfying the following criteria; cf. [taGL]*Sec. 5.
Proposition 10.20.
The functors $F$ and $F^{-1}$ are mutually inverse. Moreover, both functors send medial objects to medial objects.
Proposition 10.21.
$F(\mathsf{LQ})=\mathsf{KIR}$.
Proposition 10.22.
$F(\mathsf{LK})=\mathsf{InvRack}$.
Once we define $F$ and $F^{-1}$, these three results will more or less follow from the definitions. Together with Proposition 10.19, these results yield Theorem 10.1.
10.4.1. Construction of functors
Define $F:\mathsf{LR}\to\mathsf{Rack}$ on objects by
$(X,s,\mathtt{u})\mapsto(X,\mathtt{u}^{3}s),$
where the rack structure $\mathtt{u}^{3}s:X\to S_{X}$ is defined by $x\mapsto\mathtt{u}^{3}s_{x}$. Let $F$ preserve the underlying set maps of morphisms in $\mathsf{LR}$.
To construct an inverse functor $F^{-1}:\mathsf{Rack}\to\mathsf{LR}$, define $F^{-1}$ on objects by
$R=(X,s)\mapsto(X,\theta^{-3}_{R}s,\theta_{R}),$
where the rack structure $\theta^{-3}_{R}s:X\to S_{X}$ is defined by $x\mapsto\theta^{-3}_{R}s_{x}$. Let $F$ preserve the underlying set maps of morphisms in $\mathsf{Rack}$.
Using the inclusions $\mathtt{u}^{3}\in Z(\mathsf{GLR})$ and $\theta,\theta^{-3}\in Z(\mathsf{Rack})$, it is straightforward to verify that $F$ and $F^{-1}$ are covariant functors.
10.4.2. Proofs of Propositions 10.2010.22
Proof of Proposition 10.20.
It is clear that $FF^{-1}=\mathbf{1}_{\mathsf{Rack}}$. Using the fact that $\theta_{R}=\mathtt{u}^{-2}$ for all Legendrian racks $(R,\mathtt{u})$, verifying that $F^{-1}F=\mathbf{1}_{\mathsf{LR}}$ is straightforward. Finally, the claim that $F$ and $F^{-1}$ send medial objects to medial objects follows from the fact that, for all GL-racks $(R,\mathtt{u})$, the maps $\theta_{R}^{-3}$ and $\mathtt{u}$ are endomorphisms of $R$. ∎
Proof of Proposition 10.21.
We verify that $F(\mathsf{LQ})=\mathsf{KIR}$. If $R=(X,s)$ is a kink-involutory rack, then it is straightforward to check that the underlying rack $(X,\theta^{-1}_{R}s)$ of the Legendrian rack $F^{-1}(R)$ is a quandle.
Conversely, if $L=(X,s,\mathtt{u})$ is a Legendrian quandle, then
$\theta_{F(L)}^{2}=(\mathtt{u}^{3}\theta_{(X,s)})^{2}=\mathtt{u}^{2}=\operatorname{id}_{X},$
so $F(L)$ is kink-involutory. ∎
Proof of Proposition 10.22.
We verify that $F(\mathsf{LK})=\mathsf{InvRack}$. Let $L=(X,s,\mathtt{u})$ be a Legendrian kei. Since $\mathtt{u}$ and $s_{x}$ commute and are involutions for all $x\in X$, the rack $F(L)=(X,\mathtt{u}s)$ is involutory.
Conversely, if $R=(X,s)$ is an involutory rack, then Proposition 10.9 and Proposition 10.21 show that $F^{-1}(R)$ is a Legendrian quandle with rack structure $\theta_{R}^{-1}s$. Since $\theta^{-1}\in Z(\mathsf{Rack})$, Proposition 10.9 shows that this rack structure is involutory. Hence, $F^{-1}(R)$ is a Legendrian kei. ∎
11. Open questions
We conclude by proposing directions for future work.
Problem 11.1.
Compute the set $\operatorname{Anti}R$ for various families of racks $R$.
Problem 11.2.
Classify the group $\operatorname{Aut}R\cup\operatorname{Anti}R$ for various families of racks $R$.
Problem 11.3.
[…]
Problem 11.4.
Generalize Theorem 1.3 to $n$-fold conjugation quandles (see [matveev]*Ex. 2 in Sec. 2).
Problem 11.5.
Distinguish the symmetric quandles obtained from Theorems 1.3 and 1.4 up to isomorphism.
In light of our computational findings in Appendix B, we also pose the following two questions.
Problem 11.6.
Let $G$ be a finite nonabelian group of order $n$. Is the number of good involutions of $\operatorname{Conj}G$ completely determined by $n$, the order of the center $Z(G)$, and the class number $k(G)$ of $G$? If so, is there an exact formula to compute this number?
Problem 11.7.
In the setting of Corollary 6.12, suppose that $|Z(H)|,|X/H|\geq 2$ (so, in particular, $\operatorname{Conj}X$ is not connected). Is the upper bound on $|{\operatorname{Good}(\operatorname{Conj}X)}|$ in Corollary 6.12 always strict?
Note that if we drop either of the assumptions that $|Z(G)|,|X/H|\geq 2$, then Problem 11.7 has a negative answer; see Remark 6.13.
In light of Proposition 5.10 and Examples 9.29.5, we conclude with the following conjecture.
Conjecture 11.8.
Let $R$ be a connected quandle. If $R$ is noninvolutory, then $R$ has at most one good involution.
By Observation 2.11 and Corollary 6.11, the specialization of Conjecture 11.8 to subquandles $\operatorname{Conj}X$ of conjugation quandles $\operatorname{Conj}G$ is equivalent to the following group-theoretic conjecture.
Conjecture 11.9.
Let $G$ be a group, and let $X$ be a subset of $G$ such that $X$ is a conjugacy class of the subgroup $H:=\langle X\rangle\leq G$. If there exist distinct central elements $z_{1},z_{2}\in Z(H)$ such that $z_{1}x^{-1},z_{2}x^{-1}\in X$ for all $x\in X$, then $x^{2}\in Z(H)$ for all $x\in X$.
Appendix A GAP code for Table 4
Given an integer $n\geq 5$, the following GAP [GAP4] program produces the orders of the CNS-subquandles of $\operatorname{Conj}S_{n}$ constructed in Example 9.4. Table 4 displays the output for $5\leq n\leq 9$.
ComputeClassSizes := function(n)
local Sn, An, classes, sizes, C, rep, ord, centSn, centAn;
Sn := SymmetricGroup(n);
An := AlternatingGroup(n);
classes := ConjugacyClasses(Sn);
sizes := [];
for C in classes do
rep := Representative(C);
ord := Order(rep);
if ord > 2 then
# Use the orbit-stabilizer theorem to compare class sizes
centSn := Centralizer(Sn, rep);
centAn := Centralizer(An, rep);
# Verify that C isn’t a conjugacy class that splits in An
if SignPerm(rep) = -1 or Size(centSn) = 2 * Size(centAn)
then Add(sizes, Size(C));
fi;
fi;
od;
Sort(sizes);
return sizes;
end;
[…]
Appendix B Census of good involutions
In this section, we enumerate good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ and nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323.
Our enumerations are based on the computational data obtained from our GAP [GAP4] implementations of Algorithms 8.1 and 8.5. We provide our code and the raw data (including the explicit mappings $\rho:G\to G$) in a GitHub repository [code].
Notation.
In Tables 6 and LABEL:tab:app2, we use the following notation for groups. Let $n\geq 3$ be an integer.
  • Given a group $G$, let $k(G)$ be the number of conjugacy classes of $G$.
  • Let $S_{n}$ and $A_{n}$ denote the symmetric and alternating groups on $n$ letters, respectively.
[…]
  • Let $Q_{8}$, $QD_{16}$, and $M_{16}$ denote the quaternion group of order 888, the quasidihedral group of order 161616, and the modular maximal-cyclic group of order 161616, respectively.
In Table 6, for all nonabelian groups $G$ up to order 222222, we compare the number of good involutions of $\operatorname{Conj}G$ against the upper bound on this number from Corollary 6.12. These results motivate Problems 11.6 and 11.7. We only consider nonabelian groups because conjugation quandles of abelian groups are trivial. Cf. Table 2.
Table 6. Number of good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order 222222, compared against the upper bound from Corollary 6.12.
Group Order $||{\operatorname{Good}(\operatorname{Conj}G)}|$$|Z(G)|^{k(G)}$
$S_{3}$ 6 1 1
$D_{4}$ 8 16 32
$Q_{8}$ 8 16 32
$D_{5}$ 10 1 1
$\operatorname{Dic}_{3}$ 12 8 64
$A_{4}$ 12 1 1
$D_{6}$ 12 8 64
$D_{7}$ 14 1 1
$(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})\rtimes\mathbb{Z}/4\mathbb{Z}$ 16 2160 1048576
$\mathbb{Z}/4\mathbb{Z}\rtimes\mathbb{Z}/4\mathbb{Z}$ 16 2160 1048576
$M_{16}$ 16 2160 1048576
$D_{8}$ 16 32 128
$QD_{16}$ 16 32 128
$\operatorname{Dic}_{4}$ 16 32 128
$D_{4}\times\mathbb{Z}/2\mathbb{Z}$ 16 2160 1048576
$Q_{8}\times\mathbb{Z}/2\mathbb{Z}$ 16 2160 1048576
$(\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})\rtimes\mathbb{Z}/2\mathbb{Z}$ 16 2160 1048576
$D_{9}$ 18 1 1
$S_{3}\times\mathbb{Z}/3\mathbb{Z}$ 18 64 19683
$(\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z})\rtimes\mathbb{Z}/2\mathbb{Z}$ 18 1 1
$\operatorname{Dic}_{5}$ 20 16 256
$\mathbb{Z}/5\mathbb{Z}\rtimes\mathbb{Z}/4\mathbb{Z}$ 20 1 1
$D_{10}$ 20 16 256
$\mathbb{Z}/7\mathbb{Z}\rtimes\mathbb{Z}/3\mathbb{Z}$ 21 1 1
$D_{11}$ 22 1 1
In Table LABEL:tab:app2, for all groups $G$ of exponent greater than 222 up to order 232323, we compare the number of good involutions of $\operatorname{Core}G$ against the upper bound on this number from Corollary 7.14. We do not consider elementary abelian 222-groups because their core quandles are trivial. Cf. Table 3.
In both tables, we list groups in the order of their GAP implementations SmallGroup(n,k), where n is the order of the group and k is the index of the group in the list AllSmallGroups(n).
Table 7. Number of good involutions of nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323, compared against the upper bound from Corollary 7.14.
Group Order $||{\operatorname{Good}(\operatorname{Core}G)}|$$|T|^{|\mathcal{O}|}$
$\mathbb{Z}/3\mathbb{Z}$ 3 1 1
$\mathbb{Z}/4\mathbb{Z}$ 4 4 4
$\mathbb{Z}/5\mathbb{Z}$ 5 1 1
$S_{3}$ 6 1 1
$\mathbb{Z}/6\mathbb{Z}$ 6 2 4
$\mathbb{Z}/7\mathbb{Z}$ 7 1 1
$\mathbb{Z}/8\mathbb{Z}$ 8 4 4
$\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 8 36 256
$D_{4}$ 8 16 32
$Q_{8}$ 8 16 16
$\mathbb{Z}/9\mathbb{Z}$ 9 1 1
$\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$ 9 1 1
$D_{5}$ 10 1 1
$\mathbb{Z}/10\mathbb{Z}$ 10 2 4
$\mathbb{Z}/11\mathbb{Z}$ 11 1 1
$\operatorname{Dic}_{3}$ 12 8 8
$\mathbb{Z}/12\mathbb{Z}$ 12 4 4
$A_{4}$ 12 1 1
$D_{6}$ 12 8 64
$\mathbb{Z}/6\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 12 10 256
$\mathbb{Z}/13\mathbb{Z}$ 13 1 1
$D_{7}$ 14 1 1
$\mathbb{Z}/14\mathbb{Z}$ 14 2 4
$\mathbb{Z}/15\mathbb{Z}$ 15 1 1
$\mathbb{Z}/16\mathbb{Z}$ 16 4 4
$\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/4\mathbb{Z}$ 16 256 256
$(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})\rtimes\mathbb{Z}/4\mathbb{Z}$ 16 576 4096
$\mathbb{Z}/4\mathbb{Z}\rtimes\mathbb{Z}/4\mathbb{Z}$ 16 384 1024
$\mathbb{Z}/8\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 16 36 256
$M_{16}$ 16 16 16
$D_{8}$ 16 32 32
$QD_{16}$ 16 32 32
$\operatorname{Dic}_{4}$ 16 32 32
$\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ 16 5776 16777216
$D_{4}\times\mathbb{Z}/2\mathbb{Z}$ 16 2160 1048576
$Q_{8}\times\mathbb{Z}/2\mathbb{Z}$ 16 1296 65536
$(\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})\rtimes\mathbb{Z}/2\mathbb{Z}$ 16 256 256
$\mathbb{Z}/17\mathbb{Z}$ 17 1 1
$D_{9}$ 18 1 1
$\mathbb{Z}/18\mathbb{Z}$ 18 2 4
$S_{3}\times\mathbb{Z}/3\mathbb{Z}$ 18 1 1
$(\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z})\rtimes\mathbb{Z}/2\mathbb{Z}$ 18 1 1
$\mathbb{Z}/6\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$ 18 2 4
$\mathbb{Z}/19\mathbb{Z}$ 19 1 1
$\operatorname{Dic}_{5}$ 20 16 16
$\mathbb{Z}/20\mathbb{Z}$ 20 4 4
$\mathbb{Z}/5\mathbb{Z}\rtimes\mathbb{Z}/4\mathbb{Z}$ 20 1 1
$D_{10}$ 20 16 256
$\mathbb{Z}/10\times\mathbb{Z}/2\mathbb{Z}$ 20 10 256
$\mathbb{Z}/7\mathbb{Z}\rtimes\mathbb{Z}/3\mathbb{Z}$ 21 1 1
$\mathbb{Z}/21\mathbb{Z}$ 21 1 1
$D_{11}$ 22 1 1
$\mathbb{Z}/22\mathbb{Z}$ 22 2 4
$\mathbb{Z}/23\mathbb{Z}$ 23 1 1
Table 7. (continued)